### v1

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract
Modern deep neural networks often require distributed training with many workers due to their large size. As worker numbers increase, communication overheads become the main bottleneck in data-parallel minibatch stochastic gradient methods with per-iteration gradient synchronization. Local gradient methods like Local SGD reduce communication by only syncing after several local steps. Despite understanding their convergence in i.i.d. and heterogeneous settings and knowing the importance of batch sizes for efficiency and generalization, optimal local batch sizes are difficult to determine. We introduce adaptive batch size strategies for local gradient methods that increase batch sizes adaptively to reduce minibatch gradient variance. We provide convergence guarantees under homogeneous data conditions and support our claims with image classification experiments, demonstrating the effectiveness of our strategies in training and generalization.
1 Introduction
Recent advances in deep learning, particularly large language models (LLMs) with billions of parameters, necessitate efficient distributed training on many GPUs or TPUs. Data-parallel optimization [83] simplifies this by having each worker compute and average batch gradients from different i.i.d.  data batches, followed by a global model update. Despite its efficiency with numerous workers, this method faces challenges such as costly gradient synchronization and reduced generalization from large batches. Frequent synchronization of models or gradients among workers leads to significant communication overheads, exacerbated as worker counts and model sizes increase [71, 48, 79]. This challenge intensifies with workers distributed across nodes with limited inter-node communication speeds. To mitigate this, communication-efficient distributed optimization methods [43], such as local gradient methods [67, 82, 81, 26, 73] and their variants, have been developed. In local stochastic gradient methods (Local SGD), workers train locally and synchronize less frequently by averaging models after several local steps, effectively reducing communication overheads. In data-parallel optimization, the total batch size is the sum of all workers’ local batch sizes, increasing with more workers. Large batch sizes are preferred to fully utilize worker memory, making large-batch training the standard in large-scale model training. However, this approach often results in reduced model generalization, a problem noted by LeCun et al. [47]. Large-batch training typically creates a loss landscape with sharp minima, worsening generalization [35]. Although increasing the learning rate proportionally to batch size can mitigate the generalization gap [29, 65, 66], it cannot fully close it. Keskar et al. [35] recommended adaptive sampling to gradually increase batch sizes during training [7, 19], explored in [11, 12, 44, 59]. However, this method is still confined to single- or few-worker scenarios and does not scale well with larger models and more workers.
Motivated by the above two challenges, we answer the following question affirmatively:
Can we develop training strategies which are communication-efficient and memory-efficient with desirable generalization performance for large-scale distributed training?
Contributions.
We make the following contributions: (i) We develop adaptive batch size strategies for local gradient methods by extending the adaptive sampling methods in [7, 3]. Such strategies are for local batch sizes, i.e., batch sizes of individual workers. This is particularly relevant when the workers are heterogeneous devices with different computational speeds and memories, as well as different local objectives and datasets. The proposed strategies are communication-efficient by leveraging local gradient methods, and memory-efficient by leveraging large batch sizes at later stages of training. (ii) Under various standard assumptions, we provide convergence guarantees of the proposed approaches for smooth (strongly) convex and nonconvex objectives under a homogeneous data setting. (iii) We empirically demonstrate the efficacy of the proposed strategies with a three-layer convolutional neural network and ResNet for image classification on the MNIST and CIFAR-10 datasets, respectively, in terms of communication efficiency, memory efficiency, and model generalization performance. To grasp a thorough understanding of the strategies, we also study the effect of different hyperparameters such as the number of local steps and the one which controls the probability of increasing batch sizes.
2 Related Work
Minibatch gradient methods.
Minibatch SGD [13, 63, 4, 15] is the simplest and most widely used optimizer for large-scale machine learning [5], and is easily extendable with data parallelism [49]. The convergence of minibatch SGD and its variants like AdaGrad [18] and Adam [39] is well-understood across various settings, including convex [21], nonconvex [36], and nonsmooth [10] objectives, under different gradient variance assumptions. However, frequent gradient synchronization among workers leads to communication overheads, posing a significant challenge in distributed training.
Local gradient methods.
Local SGD [67, 83, 54], a communication-efficient variant of minibatch SGD, reduces overheads from frequent synchronization in distributed training. Recent studies [51, 17] highlight its efficacy in pre-training large language models with extensive computational resources. Optimization of Local SGD for varied data settings has shown promising convergence rates [67, 77, 76, 40, 37, 69, 74]. It includes extensions like asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75]. Federated learning [54, 9, 33, 34, 62], another model focusing on efficiency, addresses heterogeneity and partial participation. While applicable, we defer exploring federated learning’s theoretical and empirical aspects to future work.
Generalization gap in large-batch training.
Large-batch training not only faces communication overheads in minibatch SGD under distributed data parallelism but also shows poorer generalization despite better hardware efficiency. It has been successful in tasks like ImageNet classification and BERT training [80, 22, 64]. Addressing these, Lin et al. [50] introduced Post-local SGD, a hybrid of minibatch and Local SGD, where early training involves frequent synchronization of local models, followed by Local SGD. In large-scale environments, Ortiz et al. [58] noted that Post-local SGD does not consistently outperform Local SGD in generalization, and the transition from minibatch to Local SGD presents a trade-off between communication efficiency and generalization. Following these findings, Gu et al. [25, 24] analyzed Local SGD’s generalization via an SDE approximation and proposed the Quadratic Synchronization Rule (QSR) for adjusting local gradient steps during training.
Adaptive batch size strategies.
In stochastic optimization, adaptive sampling methods [7, 19, 3] adjust batch sizes based on gradient noise, enhancing training processes for smooth finite-sum problems. These methods, further explored in deep learning [11, 12, 44, 59], have not been applied to data parallelism with distributed batches. The strategies discussed here draw from batch size scheduling, a key technique in pretraining LLMs, evidenced by its use in GPT-3 [6], Nemotron-4 [60], OLMo 65B [23], and DeepSeek-V2 [14], which involve increasing batch sizes in set stages. Despite their benefits for efficiency and parallelization, these strategies are heuristic and have unclear impacts on training. This work aims to develop theoretically grounded, communication-efficient adaptive batch size strategies akin to adaptive sampling, viewed as stochastic variance-reduction methods [31, 78, 20], focusing on scalable batch size increments.
Several studies have explored adaptive batch size strategies in deep learning, including Big Batch SGD [11, 12], CABS [1], AdaBatch [16], and SimiGrad [61]. Conversely, AdaScale SGD [32] modifies learning rates for large batches instead of batch sizes. These methods, however, are often not theoretically robust, lacking solid convergence proofs or are restricted to SGD analyses under strict conditions like convexity or the Polyak–Łojasiewicz condition. Their batch size adjustment strategies usually depend on heuristic rules, such as geometric growth or decay [61], which can lead to unreliable outcomes. Crucially, these methods overlook the distributed data-parallel context, prevalent in practice, thus not addressing the needs of adaptive batch size strategies in large-scale distributed training.
Hyperparameter tuning for Local SGD.
Besides batch sizes, recent research has also focused on optimizing learning rate schedules for Local SGD, particularly in identical data settings as discussed in [2]. This optimization is crucial in federated learning with heterogeneous data, where local objectives differ significantly in geometry and smoothness, necessitating locally adaptive learning rate schedules [56, 38]. Our adaptive local batch size strategies complement these efforts.
3 Preliminaries
Notation.
[…]
3.1 Problem Formulation
We consider a distributed training setting with $M$ workers, each with a possibly heterogeneous (i.e. non-i.i.d. ) underlying true data distribution $\mathbb{P}_{m}$ ($m\in\llbracket M\rrbracket$). The true data distribution $\mathbb{P}_{m}$ is usually unknown, but can be approximated by its empirical distribution ^$\widehat{\mathbb{P}}_{m}=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}\updelta_{\xi_{i}^{m}}$, where $\mathcal{D}_{m}\coloneqq\{\xi_{i}^{m}\}_{i\in\llbracket n_{m}\rrbracket}$ is the set of $n_{m}$ data samples of worker $m$ and $\xi_{i}^{m}\in\mathcal{Z}_{m}\subseteq\mathbb{R}^{p}$ for $i\in\llbracket n_{m}\rrbracket$. The goal is to find an approximate minimizer of the global objective $F$ which is the average of the local objectives $F_{m}$:
$\operatorname*{minimize}_{x\in\mathbb{R}^{d}}\ F(x)\coloneqq\frac{1}{M}\sum_{m=1}^{M}F_{m}(x),$
where $F_{m}(x)\coloneqq\mathbb{E}_{\xi\sim\widehat{\mathbb{P}}_{m}}[f_{m}(x;\xi)]=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}f_{m}(x;\xi_{i}^{m})$ is the local objective for worker $m$.
At each iteration $k\in\mathbb{N}$, since the number of samples of each worker $n_{m}$ could be large, the per-worker gradient $\nabla F_{m}$ cannot be computed. It can however be approximated by its minibatch counterpart
$(\forall x\in\mathbb{R}^{d})\quad\nabla F_{\mathcal{B}_{k}^{m}}(x)\coloneqq\frac{1}{b_{k}^{m}}\sum_{i\in\mathcal{B}_{k}^{m}}\nabla f_{m}(x;\xi_{i}^{m}),$ (1)
where $\mathcal{B}_{k}^{m}$ is the local batch of worker $m$ at iteration $k$ and $b_{k}^{m}\coloneqq|\mathcal{B}_{k}^{m}|$ is the corresponding local batch size, assuming that $f_{m}(\cdot;\xi^{m})$ is continuously differentiable for any $\xi^{m}\in\mathcal{Z}_{m}$.
3.2 Minibatch and Local Stochastic Gradient Methods
We present both the formulations of minibatch SGD under data parallelism and Local SGD.
Minibatch SGD.
Minibatch SGD under data parallelism is a fully synchronized stochastic gradient method. In particular, at each iteration $k$, after each worker $m$ has computed its local batch gradient $\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$, the global batch gradient $\nabla F_{\mathcal{B}_{k}}(x_{k})$ is computed by averaging all local batch gradients by $\nabla F_{\mathcal{B}_{k}}(x_{k})=\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$ with a global batch size $b_{k}\coloneqq|\mathcal{B}_{k}|=\sum_{m=1}^{M}b_{k}^{m}=Mb_{k}^{1}$, assuming that all local batch sizes are equal. This is followed by the global model update, performed via $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$, where $\alpha_{k}$ is the learning rate of iteration $k$. The same mechanism also applies to other optimizers such as Adam [39] and AdaGrad [18].
Local SGD.
As opposed to minibatch SGD, Local SGD [67] reduces the communication frequency. Communication for model averaging (through all-reduce operations) is performed every $H$ local gradient steps, Local SGD takes the following updates:
$(\forall m\in\llbracket M\rrbracket)(\forall k\in\mathbb{N})\quad x_{k+1}^{m}=\begin{cases}\frac{1}{M}\sum_{m=1}^{M}\left(x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right)&\text{if }H\text{ divides }k+1,\\ x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})&\text{otherwise}.\end{cases}$ (2)
We also consider the following alternative representation that explicitly separates the numbers of local gradient steps $H$ and communication rounds $K$. In particular, for $M$ parallel workers, $K$ rounds of communication, $H$ local gradient steps per round, Local SGD updates can also be expressed as
$(\forall m\in\llbracket M\rrbracket)(\forall k\in\llbracket 0,K-1\rrbracket)\;\begin{cases}x_{k,h+1}^{m}=x_{k,h}^{m}-\alpha_{k,h}\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})&\text{for }h\in\llbracket 0,H-1\rrbracket,\\ x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}.\end{cases}$ (3)
4 Adaptive Local Batch Size Strategies for Local Gradient Methods
The proposed adaptive local batch size strategies are based on the adaptive sampling method [7] originally developed for unconstrained stochastic optimization problems using minibatch SGD. We outline its mechanism and its extension to Local SGD below.
4.1 Adaptive Sampling Methods
Introduced in [7], the adaptive sampling method for minibatch SGD was designed for a single worker ($M=1$) with the objective $F(x)=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i})$ over samples $\mathcal{D}=\{\xi_{i}\}_{i=1}^{n}$. This method relies on a key stochastic optimization principle: when the sample size $n$ is large, the batch gradient $\nabla F_{\mathcal{B}_{k}}$ approximates the full gradient $\nabla F$. If $f(\cdot;\xi)$ is continuously differentiable and convex for any $\xi\in\mathcal{Z}\subseteq\mathbb{R}^{p}$, then $-\nabla F_{\mathcal{B}_{k}}(x_{k})$ serves as a descent direction for $F$ at $x_{k}\in\mathbb{R}^{d}$, provided there exists $\eta\in\left(0,1\right)$ such that
$\delta_{\mathcal{B}_{k}}(x_{k})\coloneqq\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|\leqslant\eta\|\nabla F(x_{k})\|.$ (4)
Note that $\nabla F(x_{k})$ is not available in minibatch SGD, but $\delta_{\mathcal{B}_{k}}(x_{k})$ can be approximated by $\widehat{\delta}_{\mathcal{B}_{k}}(x_{k})^{2}\coloneqq\frac{1}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))\cdot\frac{n-b_{k}}{n-1}$, where, for any vector-valued function $h\colon\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}^{d}$, the batch variance is defined by
$\mathrm{Var}_{i\in\mathcal{B}_{k}}(h(x_{k};\xi_{i}))\coloneqq\frac{1}{b_{k}-1}\sum_{i\in\mathcal{B}_{k}}\left\lVert h(x_{k};\xi_{i})-\mathbb{E}_{i\in\mathcal{B}_{k}}[h(x_{k};\xi_{i})]\right\rVert^{2},$ (5)
with $\mathbb{E}_{i\in\mathcal{B}_{k}}[h(x_{k};\xi_{i})]\coloneqq\frac{1}{b_{k}}\sum_{i\in\mathcal{B}_{k}}h(x_{k};\xi_{i})$. Hence, as $n\to\infty$, condition (4) can be approximated as
$\frac{1}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))\leqslant\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}.$ (6)
In practice, for each iteration $k\in\mathbb{N}$, the dynamic sample gradient algorithm [7] performs the following minibatch SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with learning rate $\alpha_{k}>0$, and then checks the following (approximate) norm test condition (6), which can also be viewed as an approximation of the exact variance norm test in the stochastic setting:
$\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|^{2}\right]\leqslant\eta^{2}\|\nabla F(x_{k})\|^{2},$ (7)
i.e., condition (4) holds in expectation. Here we abbreviate the conditional expectation on $\mathcal{F}_{k}$ (i.e., the filtration $\mathcal{F}_{k}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$. If condition (6) is violated, the batch size of the next iteration $b_{k+1}$ is increased, determined by the formula:
$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\rceil,$
and the next batch $\mathcal{B}_{k+1}$ is sampled with this larger batch size accordingly. Otherwise, the next batch size remains the same as the current one. The value of the pre-specified constant $\eta$ controls the probability of obtaining a descent direction, thereby affecting the probability of increasing the batch size as well as the magnitude of the next batch size.
Condition (7) is more relaxed than the typical assumption of uniformly bounded gradient variance for minibatch stochastic gradient methods: for any i.i.d.  batch $\mathcal{B}\subset\llbracket n\rrbracket$ and some positive constant $\sigma>0$, $\mathbb{E}_{\mathcal{B}}\left[\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|^{2}\right]\leqslant\sigma^{2}$ for all $x\in\mathbb{R}^{d}$, which is usually unverifiable in practice. The adaptive batch size schedule using the approximate norm test may cause rapid batch size increases, undermining its purpose. To moderate these schedules, Bollapragada et al. [3] introduced the (augmented) inner product test, which manages the variance between the batch and full gradient’s inner product. Given the preference for large batch sizes for memory efficiency, controlled via $\ eta$, we focus on the norm test and defer exploring local variants of the (augmented) inner product test to future work.
4.2 Extensions to Local Gradient Methods
We extend the above adaptive sampling methods beyond the single worker setting with data parallelism using multiple workers, more specifically for Local SGD. We have to invoke a local variant of the norm test (7), called the (exact variance) local norm test, at every iteration $k\in\mathbb{N}$, using formulation (2): for some sequence of positive constants $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$,
$(\forall m\in\llbracket M\rrbracket)\quad\mathbb{E}_{\mathcal{F}_{k}^{m}}\left[\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})-\nabla F_{m}(x_{k}^{m})\|^{2}\right]\leqslant\eta_{m}^{2}\left\lVert\nabla F_{m}(x_{k}^{m})\right\rVert^{2},$ (8)
where $\mathbb{E}_{\mathcal{F}_{k}^{m}}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}^{m}]$ denotes the conditional expectation on $\mathcal{F}_{k}^{m}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0}^{m},\ldots,\mathcal{B}_{k-1}^{m}\})$. Again, $\nabla F_{m}$ is in general unavailable since the sample size $n_{m}$ could be large. The above local norm test is approximately enforced by the following (approximate) local norm test:
$(\forall m\in\llbracket M\rrbracket)\quad\frac{1}{b_{k}^{m}}\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}\left(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m})\right)\leqslant\eta_{m}^{2}\left\lVert\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right\rVert^{2}.$ (9)
Should the condition (9) not be met, the subsequent local batch sizes are determined by the formula:
$(\forall m\in\llbracket M\rrbracket)\quad b_{k+1}^{m}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m}))}{\eta_{m}^{2}\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\|^{2}}\right\rceil,$ (10)
and the local batches $\mathcal{B}_{k+1}^{m}$ are augmented accordingly. In a practical implementation, for simplicity, we can set $\eta_{m}\equiv\eta$ for all $m\in\llbracket M\rrbracket$ and choose the next local batch sizes to be
$(\forall m\in\llbracket M\rrbracket)\quad b_{k+1}^{m}\equiv b_{k+1}\coloneqq\max_{m\in\llbracket M\rrbracket}\ \left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\|^{2}}\right\rceil,$ (11)
which is particularly useful when homogeneous workers and identical data on the workers are used, so that there would not be discrepancy of training time for the local steps due to different local batch sizes and undesirable stragglers before model averaging.
The proposed adaptive batch size strategies extend to local variants of minibatch stochastic gradient optimizers beyond SGD [63], including momentum SGD [70], AdaGrad [18], Adam [39], and AdamW [52]. The algorithm is detailed in Algorithm 1.
In a data parallel setting, adaptive global batch size strategies can be developed using minibatch SGD, leading to increased communication overheads from aggregating local batch gradient variances among workers to compute the quantities in (6).
Algorithm 1 Adaptive Local Batch Size Schedules for Local Gradient Methods
no. of workers $M\in\mathbb{N}^{*}$, no. of communication steps $K\in\mathbb{N}^{*}$, no. of local steps $H\in\mathbb{N}^{*}$, no. of training samples $N\in\mathbb{N}^{*}$, global step counter $k=0$, processed samples counter $B=0$, training sets $(\mathcal{D}_{m})_{m\in\llbracket M\rrbracket}=(\{\xi_{i}^{m}\}_{i=1}^{n_{m}})_{m\in\llbracket M\rrbracket}$, initial model parameters $x_{0,0}^{m}=x_{0}\in\mathbb{R}^{d}$for all $m\in\llbracket M\rrbracket$, initial local batch sizes $(b_{0}^{m})_{m\in\llbracket M\rrbracket}$, learning rate schedules $(\alpha_{k,h})_{k\in\llbracket 0,K-1\rrbracket,h\in\llbracket 0,H-1\rrbracket}$, $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$$B<N$worker $m=1,\ldots,M$in parallel local step $h=0,\ldots,H-1$
Sample the local batch $\mathcal{B}_{k,h}^{m}$ of size $b_{k,h}^{m}\coloneqq|\mathcal{B}_{k,h}^{m}|$ uniformly from $\mathcal{D}_{m}$
Compute per-sample gradients $\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$ for each $i\in\mathcal{B}_{k,h}^{m}$
Compute the local batch gradient $\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})=\frac{1}{b_{k,h}^{m}}\sum_{i\in\mathcal{B}_{k,h}^{m}}\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$
Compute $\mathrm{Var}_{i\in\mathcal{B}_{k,h}^{m}}(\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m}))$ by applying (5)
Compute the local norm test statistic $\mathsf{T}_{k,h}^{m}$ defined by (10)
Determine the next local batch sizes: $b_{k,h+1}^{m}=\max\{\mathsf{T}_{k,h}^{m},b_{k,h}^{m}\}$
[…]
Update the global model: $x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}$ for each $m\in\llbracket M\rrbracket$triangleright all-reduce
$B\leftarrow B+\sum_{m=1}^{M}\sum_{h=0}^{H-1}b_{k,h}^{m}$
[…]
5 Convergence Analysis
Our analysis focuses on the i.i.d.  setting as shown in (1). We set $\eta_{m}\equiv\eta\in(0,1)$ for all $m\in\llbracket M\rrbracket$, typical in datacenter environments for distributed training [13], where workers access the entire data set. The analysis of heterogeneous data settings is deferred to future work. Our convergence analysis, following [69], does not assume bounded gradient variance across $\mathbb{R}^{d}$. Instead, our adaptive local batch size strategies ensure the expected strong growth condition [36, 5] at each local iterate $x_{k}^{m}$ for all workers $m\in\llbracket M\rrbracket$. This approach relaxes the standard assumptions of uniform boundedness and expected strong growth conditions [72], which are challenging to verify except in the interpolation regime [53].
We might need to invoke the following assumptions on the objective function $F$.
Assumption 1 ($L$-Lipschitz smoothness).
The function $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable, bounded below by $F^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}F(x)=F(x^{\star})\in\mathbb{R}$ with $x^{\star}=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}F(x)$, and is $L$-Lipschitz smooth for some $L>0$, i.e., for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $\left\lVert\nabla F(x)-\nabla F(y)\right\rVert\leqslant L\left\lVert x-y\right\rVert$.
Assumption 2 ($\mu$-strong convexity).
The function $F$ is $\mu$-strongly convex for some $\mu\geqslant 0$, i.e., for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $F(x)-F(y)+\frac{\mu}{2}\left\lVert x-y\right\rVert^{2}\leqslant\langle\nabla F(x),x-y\rangle$. If $\mu=0$, we say that $F$ is convex.
We derive the following convergence results for the proposed adaptive batch size strategies as follows. Their proofs are deferred to Appendix A.
Theorem 1 (Strongly convex; $\mu>0$).
Suppose that Assumptions 1 and 2 hold with $\mu>0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$ and a constant learning rate $\alpha_{k}\equiv\alpha$. Then, for some $\alpha\leqslant 1/(10L(HM+\eta^{2}))$, we have
$\mathbb{E}[F(x_{\mathsf{out}})]-F^{\star}=\widetilde{\mathscr{O}}\left(L(HM+\eta^{2})\exp\left\{-\frac{\mu K}{10L(HM+\eta^{2})}\right\}\left\lVert x_{0}-x^{\star}\right\rVert^{2}\right),$ (12)
where $x_{\mathsf{out}}\in\{x_{k}^{m}\}_{k\in\llbracket 0,K-1\rrbracket,m\in\llbracket M\rrbracket}$ is chosen as $x_{k}^{m}$ with probability proportional to $(1-\mu\alpha/2)^{-k}$ uniformly on $m\in\llbracket M\rrbracket$ and $\widetilde{\mathscr{O}}$ hides logarithmic factors.
Theorem 2 (Convex; $\mu=0$).
Suppose that Assumptions 1 and 2 hold with $\mu=0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$ and a constant learning rate $\alpha_{k}\equiv\alpha$. Then, for some $\alpha\leqslant 1/(10L(HM+\eta^{2}))$, we have
$\mathbb{E}[F(x_{\mathsf{out}})]-F^{\star}=\mathscr{O}\left(\frac{L(HM+\eta^{2})}{K}\left\lVert x_{0}-x^{\star}\right\rVert^{2}\right),$ (13)
where $x_{\mathsf{out}}$ is chosen uniformly at random from $\{x_{k}^{m}\}_{k\in\llbracket 0,K-1\rrbracket,m\in\llbracket M\rrbracket}$.
Theorem 3 (Nonconvex).
Suppose that Assumption 1 holds. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$ and a constant learning rate $\alpha_{k}\equiv\alpha$. Then, for some $\alpha\leqslant 1/(10L(HM+\eta^{2}))$, we have
$\mathbb{E}\left\lVert\nabla F(x_{\mathsf{out}})\right\rVert^{2}=\mathscr{O}\left(\frac{L(HM+\eta^{2})}{K}(F(x_{0})-F^{\star})\right),$ (14)
where $x_{\mathsf{out}}$ is chosen uniformly at random from $\{x_{k}^{m}\}_{k\in\llbracket 0,K-1\rrbracket,m\in\llbracket M\rrbracket}$.
The theorems confirm that Local SGD, using the local norm test with exact variance, achieves linear convergence rates for both convex and nonconvex objectives. This scaling is proportional to the number of local gradient steps $H$ and workers $M$, akin to Local SGD under constant batch sizes and bounded gradient variance [69].
Remark 1 (Choice of $\eta$).
Convergence rates depend quadratically on $\eta$, which controls the likelihood and size of increased local batch sizes. Choosing smaller $\eta$ values accelerates convergence but causes quick batch size escalation, undermining adaptive batch sizing and widening the generalization gap.
6 Numerical Experiments
We evaluate the efficiency of the training (total steps or wall-clock time) and the generalization performance of our strategies through image classification tasks. We use a three-layer convolutional neural network on the MNIST dataset [46] and a ResNet-50 [28] on the CIFAR-10 dataset [42], with four workers ($M=4$ GPUs). The models are trained using Local Adam and Local SGD with momentum (SHB), synchronizing only the models every $H$ steps. Further details and results are in Appendix B.
6.1 Three-layer CNN on MNIST
We demonstrate our strategies using a three-layer CNN for MNIST classification. The training involves 12 million samples (200 epochs) with global and local batch sizes set to 60,000 and 15,000, respectively. This task uses a fixed learning rate of 0.0008 without any rate schedules. Results are in Table 1.
Table 1: Adaptive local batch size strategies for three-layer CNN on MNIST; no. of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 1456 5.01 2048 99.10 1464 9.44 2048 99.06 1464 25.58 2048 99.08
Constant 352 6.84 8192 98.80 364 8.19 8192 98.85 366 12.37 8192 98.94
Constant 192 10.59 15000 98.40 196 11.41 15000 98.43 199 13.88 15000 98.54
$\eta=0.25$ 213 11.35 13414 98.51 221 11.94 13427 98.55 223 14.50 13467 98.66
$\eta=0.375$ 224 11.85 13077 98.61 228 12.16 13031 98.58 228 14.86 13078 98.71
$\eta=0.50$ 229 11.85 12615 98.58 235 11.73 12648 98.60 235 14.42 12748 98.72
Figure 1: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies ($H=16$).
Values of $\eta$.
Table 1 shows that adaptive local batch size strategies can match or exceed the validation accuracy of constant batch size strategies. Specifically, with 16 local gradient steps and $\eta=0.25$, the adaptive strategy achieves an average batch size similar to a constant batch size of 15,000 (see Figure 1), but with superior accuracy and 21 additional steps. This confirms that adaptive strategies can mitigate the generalization gap in large-batch training. Furthermore, Figure 2 indicates that these strategies reach a maximum local batch size of 15,000 later in the training, enhancing hardware utilization and efficiency. Similar trends are noted for $H=4$ and $H=1$.
Figure 2: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
Local gradient steps $H$.
Smaller $H$ increases training time due to more frequent communication, and contrary to expectations, does not always enhance generalization. As Table 1 demonstrates, with a local batch size of 2048, $H=1$ and $H=4$ have lower accuracy compared to $H=16$. This pattern is consistent with adaptive batch sizes, where $H=4$ underperforms $H=16$ at $\eta=0.375$. Moreover, while $H=1$ leads to longer training times, it achieves higher accuracy.
6.2 ResNet-50 on CIFAR-10
We scale our experiment to ResNet-50 for CIFAR-10, using a training budget of 12 million samples (200 epochs) and setting maximum global and local batch sizes of 60,000 and 15,000, respectively. Our learning rate schedules include a 10% linear warmup and cosine decay, with peak and minimum rates of 0.5 and 0.05, respectively.
Table 2: Adaptive local batch size strategies for ResNet-50 on CIFAR-10; no. of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 2432 27.16 1024 92.31 2440 34.00 1024 92.65 2441 60.51 1024 91.11
Constant 608 34.99 4096 63.50 608 37.10 4096 57.04 610 46.39 4096 40.40
Constant 304 34.31 8192 12.51 304 35.34 8192 16.61 305 40.01 8192 14.78
$\eta=0.8$ 576 90.73 4306 81.95 416 138.39 6001 75.09 241 50.67 10380 25.32
$\eta=0.85$ 512 92.41 4900 81.95 912 141.51 2742 86.89 433 144.75 5767 72.66
$\eta=0.9$ 2816 96.98 885 93.59 1982 126.38 1261 92.53 3940 111.92 634 93.75
Values of $\eta$.
Table 3 shows that adaptive local batch size strategies can achieve comparable or superior validation accuracy to constant batch sizes with similar averages. For instance, with $H=16$ and $\eta=0.80$, the adaptive strategy reaches an average batch size akin to a constant batch size of 4096 but with nearly 50% higher accuracy and 32 fewer steps, illustrating the effectiveness of adaptive strategies in bridging the generalization gap in large-batch training.
Figure 3: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$).
Local gradient steps $H$.
We first examine constant batch sizes (Table 3 and Figure 4). Local gradient methods, as noted in [50, 58, 24, 25], may improve generalization compared to synchronized methods ($H=1$), although we employ SHB in synchronized models instead of standard SGD. For adaptive local batch sizes, smaller $H$ values increase batch sizes more rapidly (Table 3), supporting our theory that a smaller $\eta$ should be used with a larger $H$ to maintain convergence rates of $\mathscr{O}(HM+\eta^{2})$. Refer to Figures 3 and B.2 for further illustrations.
Figure 4: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
Number of workers $M$.
We conducted additional experiments with $M=8$. Compared to the 4-worker scenario, increasing the number of workers reduces training time but lowers average accuracy, indicating poorer generalization. Nonetheless, the observations regarding local gradient steps $H$ and values of $\eta$ remain consistent with the 4-worker case.
Figure 5: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$) with 8 workers.
Table 3: Adaptive local batch size strategies for ResNet-50 on CIFAR-10 with 8 workers
Schedule $H=16$ $H=4$ $H=1$
steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
Constant 2432 12.06 512 92.06 2440 15.45 512 92.87 2441 29.09 512 92.41
Constant 1216 11.19 1024 88.88 1220 12.93 1024 87.87 1220 19.72 1024 86.80
Constant 608 10.72 2048 66.94 608 11.59 2048 71.32 610 15.03 2048 59.57
$\eta=0.8$ 213 11.55 5666 17.88 216 11.82 5655 15.98 218 12.29 5650 16.84
$\eta=0.85$ 220 12.22 5610 19.47 222 12.30 5570 22.06 222 12.66 5569 21.99
$\eta=0.9$ 261 15.25 5236 28.78 277 16.92 5052 33.07 278 18.26 4959 36.83
7 Concluding Remarks
This work introduces adaptive local batch size strategies for local gradient methods in data parallelism, focusing on reducing communication overheads by increasing local batch sizes independently. Although our analysis and experiments are limited to identical data settings, the algorithm is applicable to heterogeneous scenarios. Future work will explore convergence in diverse settings and federated learning [54, 33], including personalized approaches [27]. Extensions to asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75] are also promising for improving communication efficiency and addressing worker heterogeneity.
Acknowledgments
The research of Han Liu is supported by NIH R01LM01372201, NSF RI 1840857, NSF TRIPOD 1740735, NSF DMS1454377-CAREER, NSF IIS 1546482, along with an Alfred P. Sloan Fellowship. The research of Mladen Kolar is supported in part by NSF ECCS-2216912. This research is supported in part through the computational resources and staff contributions from the Data Science Institute at the University of Chicago, through its AI + Science Research Funding Initiatives.
Appendix
Appendix A Proofs of Main Text
We present the omitted proofs of the main text, mainly by simplifying the technique in [69].
Notation.
Let us define
$(\forall k\in\mathbb{N})\quad\overline{x}_{k+1}\coloneqq\overline{x}_{k}-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}g_{k}^{m},$ (A.1)
where $g_{k}^{m}\coloneqq\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$ and $\overline{x}_{0}\coloneqq x_{0}$.
Let use recall that $\mathbb{E}_{k}\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$, $\mathbb{E}_{\mathcal{F}_{k}^{m}}\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}^{m}]$, and $\mathbb{E}$ denotes the unconditional expectation.
We also define the notion of slowly decreasing learning rates, called $\tau$-slow sequences.
Definition A.1 ($\tau$-slow sequence; Definition 10 in in [69]).
A sequence $(a_{k})_{k\in\mathbb{N}}\subset(0,\infty)$ with positive values is $\tau$-slow decreasing for $\tau\geqslant 1$ if
$(\forall k\in\mathbb{N})\quad a_{k+1}\leqslant a_{k}\qquad\text{and}\qquad a_{k+1}\left(1+\frac{1}{2\tau}\right)\geqslant a_{k}.$
The sequence $(a_{k})_{k\in\mathbb{N}}$ is $\tau$-slow increasing if $(1/a_{k})_{k\in\mathbb{N}}$ is $\tau$-slow decreasing.
A.1 Technical Lemmas
We first state without proofs some preparatory lemmas, which are either standard results or whose proofs are be found in the cited work.
Lemma A.1.
For any $M$ vectors $v_{1},\ldots,v_{M}\in\mathbb{R}^{d}$, we have
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert\leqslant\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert.$ (A.2)
Lemma A.2.
For any $M$ vectors $v_{1},\ldots,v_{M}\in\mathbb{R}^{d}$, we have
$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert^{2}\leqslant M\sum_{m=1}^{M}\left\lVert w_{m}\right\rVert^{2}.$ (A.3)
Lemma A.3.
For any two vectors $x,y\in\mathbb{R}^{d}$ and any positive number $\rho>0$, we have
$\left\lVert x+y\right\rVert^{2}\leqslant(1+\rho)\left\lVert a\right\rVert^{2}+(1+1/\rho)\left\lVert b\right\rVert^{2}.$ (A.4)
Lemma A.4 (Cauchy–Schwarz inequality).
For any two vectors $x,y\in\mathbb{R}^{d}$ and any positive number $\rho>0$, we have
$2\langle x,y\rangle\leqslant\rho\left\lVert a\right\rVert^{2}+\rho^{-1}\left\lVert b\right\rVert^{2}.$ (A.5)
Lemma A.5.
If $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is $L$-Lipschitz smooth for some $L>0$ (Assumption 1), then we have
$(\forall x\in\mathbb{R}^{d})\quad\left\lVert\nabla F(x)\right\rVert^{2}\leqslant 2L(F(x)-F^{\star}).$ (A.6)
Lemma A.6 (Lemma 2 in [68]; Lemma 13 in [69]).
For any nonnegative sequence $(u_{k})_{k\in\mathbb{N}}$ and parameters $\kappa\geqslant\chi>0$, $\sigma\geqslant 0$ and $K\in\mathbb{N}$, there exists a constant $\alpha\leqslant 1/\kappa$ such that for constant learning rates $(\alpha_{k}\equiv\alpha)_{k\in\mathbb{N}}$ and weights $w_{k}\coloneqq(1-\chi\alpha)^{-k-1}$, we have
$\Psi_{K}\coloneqq\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{(1-\chi\alpha_{k})u_{k}}{\alpha_{k}}-\frac{u_{k+1}}{\alpha_{k}}+\sigma\alpha_{k}\right)=\widetilde{\mathscr{O}}\left(\kappa u_{0}\mathrm{e}^{-\chi K/\kappa}+\frac{\sigma}{\chi K}\right),$
where $W_{K}\coloneqq\sum_{k=1}^{K}w_{k}$.
Lemma A.7 (Lemma 14 in [69]).
For any nonnegative sequence $(u_{k})_{k\in\mathbb{N}}$ and parameters $\kappa\geqslant 0$, $\sigma\geqslant 0$ and $K\in\mathbb{N}$, there exists a constant $\alpha\leqslant 1/\kappa$ such that for constant learning rates $(\alpha_{k}\equiv\alpha)_{k\in\mathbb{N}}$, we have
$\Psi_{K}\coloneqq\frac{1}{K+1}\sum_{k=0}^{K}\left(\frac{u_{k}}{\alpha_{k}}-\frac{u_{k+1}}{\alpha_{k}}+\sigma\alpha_{k}\right)\leqslant\frac{\kappa u_{0}}{K+1}+2\sqrt{\frac{\sigma u_{0}}{K+1}}.$
A.2 Proof of Theorems 1, 2 and 3
We first prove a descent lemma for ($\mu$-strongly) convex objectives.
Lemma A.8 (Descent lemma for ($\mu$-strongly) convex objectives).
Suppose that Assumptions 1 and 2 hold with $\mu>0$. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
$\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\leqslant\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{\alpha_{k}}{2M}\sum_{m=1}^{M}\mathbb{E}[F(x_{k}^{m})-F^{\star}]+\frac{3L\alpha_{k}}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
Proof.
The proof is similar to that of Lemma 25 in [69]. We state it here for convenience and completeness. Expanding we have
$\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}=\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle g_{k}^{m},x_{k}^{m}-x^{\star}\rangle+\frac{\alpha_{k}^{2}}{M^{2}}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}+\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle g_{k}^{m},x_{k}^{m}-\overline{x}_{k}\rangle.$
Taking expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\\ \leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle+\frac{\alpha_{k}^{2}}{M^{2}}\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}+\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-\overline{x}_{k}\rangle.$
Note that $\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$ can be bounded by
$\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$ $=\left\lVert\sum_{m=1}^{M}\nabla F(x_{k}^{m})\right\rVert^{2}+\sum_{m=1}^{M}\mathbb{E}\left\lVert g_{k}^{m}-\nabla F(x_{k}^{m})\right\rVert^{2}$
$\leqslant M\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\eta^{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.3) and (8) (A.7)
[…]
On the other hand, by Assumption 2, we also have
$-2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle\leqslant-\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}-2\left(F(x_{k}^{m})-F^{\star}\right).$ (A.8)
By Cauchy–Schwarz inequality (A.5) with $\rho=2L$, we have
$2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-\overline{x}_{k}\rangle$ $\leqslant 2L\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+\frac{1}{2L}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$
[…]
By (A.4), we also have
$-\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}\leqslant-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$ (A.10)
Thus, plugging in (A.6), (A.9) and (A.10), we obtain
$\mathbb{E}_{k}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}$ $\leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle+\frac{2L\alpha_{k}^{2}(M+\eta^{2})}{M^{2}}\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)$
$\qquad+\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-\overline{x}_{k}\rangle$
$\leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\left[\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+2\left(F(x_{k}^{m})-F^{\star}\right)\right]$
$\qquad+\frac{2L(M+\eta^{2})\alpha_{k}^{2}}{M^{2}}\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\left[2L\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+F(x_{k}^{m})-F^{\star}\right]$
$\leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\frac{\mu\alpha_{k}}{M}\sum_{m=1}^{M}\left[-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\right]$
$\qquad-\frac{\alpha_{k}}{M}\left(1-\frac{2L(M+\eta^{2})\alpha_{k}}{M}\right)\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)+\frac{2L\alpha_{k}}{M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant\left(1-\frac{\mu\alpha_{k}}{2}\right)\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{\alpha_{k}}{2M}\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)+\frac{3L\alpha_{k}}{M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2},$
since $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ and $\mu\leqslant L$ if $F$ is both $\mu$-strongly convex and $L$-Lipschitz smooth. Taking total expectation gives the desired result. ∎
We also prove a descent lemma for general nonconvex objectives.
Lemma A.9 (Descent lemma for nonconvex objectives).
Suppose that Assumption 1 holds. Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$. If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
$\mathbb{E}[F(\overline{x}_{k+1})]\leqslant\mathbb{E}[F(\overline{x}_{k})]-\frac{\alpha_{k}}{4M}\sum_{m=1}^{M}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{L^{2}\alpha_{k}}{2M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
Proof.
The proof is similar to that of Lemma 26 in [69]. We state it here for convenience and completeness.
By Assumption 1, we have
$F(\overline{x}_{k+1})\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),g_{k}^{m}\rangle+\frac{L\alpha_{k}^{2}}{2M^{2}}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}.$
Taking expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}[F(\overline{x}_{k+1})]$ $\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}}{2M^{2}}\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$
$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.7)
$=F(\overline{x}_{k})-\left(\frac{\alpha_{k}}{M}-\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$
$\qquad+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle.$
By Cauchy–Schwarz inequality (A.5) with $\rho=1$ and Assumption 1, we have
$\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle$ $\leqslant\frac{1}{2}\sum_{m=1}^{M}\left[\left\lVert\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k})\right\rVert{\nabla F(x_{k}^{m})}^{2}+\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}\right]$
$\leqslant\frac{L^{2}}{2}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}+\frac{1}{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}.$
Plugging this in, we obtain
$\mathbb{E}_{k}[F(\overline{x}_{k+1})]$ $=F(\overline{x}_{k})-\frac{\alpha_{k}}{2M}\left(1-\frac{L(M+\eta^{2})\alpha_{k}}{M}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{L^{2}\alpha_{k}}{2M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{4M}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{L^{2}\alpha_{k}}{2M}\sum_{m=1}^{M}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2},$
since $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$. Taking total expectation gives the desired result. ∎
We then derive an upper bound for $\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$.
Lemma A.10.
Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$, and $(\overline{x}_{k})_{k\in\mathbb{N}}$ be the sequence defined in (A.1). If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$ and $(\alpha_{k}^{2})_{k\in\mathbb{N}}$ is $H$-slow decreasing, then we have
$\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{10LHM}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.11)
Moreover, for any $H$-slow increasing nonnegative sequence $(w_{k})_{k\in\mathbb{N}}$ we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$
Proof.
We first consider
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$ $=\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{\lfloor K/H\rfloor H}-(\overline{x}_{k}-\overline{x}_{\lfloor K/H\rfloor H})\right\rVert^{2}$
$\leqslant\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{\lfloor K/H\rfloor H}\right\rVert^{2}$
$\leqslant\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert\sum_{h=0}^{H-1}\alpha_{k-h}\left(\nabla F(x_{k-h}^{m})+g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right)\right\rVert^{2}$
$\leqslant\frac{3H}{2M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
$\qquad\qquad+\frac{3}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right\rVert^{2}$ by (A.4) and (A.3)
$\leqslant\frac{3(H+\eta^{2})}{2K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$
The first inequality follows from $\mathbb{E}\left\lVert X-\mathbb{E}X\right\rVert^{2}\leqslant\mathbb{E}\left\lVert X\right\rVert^{2}$ for any random variable $X$ in $\mathbb{R}^{d}$. Since $(\alpha_{k}^{2})_{k\in\mathbb{N}}$ is $H$-slow decreasing, for $h\leqslant H$, we have $\alpha_{k-h}^{2}\leqslant\alpha_{k}^{2}(1+1/(2H))^{H}\leqslant\alpha_{k}^{2}\mathrm{e}^{H/(2H)}\leqslant 2\alpha_{k}^{2}$ since $1+x\leqslant\mathrm{e}^{x}$ for any $x\in\mathbb{R}$. Then, we can simply the inequality:
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{3(H+\eta^{2})\alpha_{k}^{2}}{K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.12)
Since $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$, we have
$3L\cdot 3(H+\eta^{2})\alpha_{k}^{2}\leqslant\frac{1}{10L(HM+\eta^{2})}\leqslant\frac{1}{10LHM}.$
Since $(w_{k})_{k\in\mathbb{N}}$ is $H$-slow increasing, we have $w_{k}\leqslant w_{k-h}(1+2/(2H))^{h}\leqslant w_{k-h}(1+2/(2H))^{H}\leqslant w_{k-h}\sqrt{\mathrm{e}}\leqslant 2w_{k-h}$ for each $h\in\llbracket 0,H\rrbracket$. Therefore, we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$ $\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}\frac{w_{k}}{2H}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}\frac{1}{H}\sum_{h=0}^{H-1}w_{k-h}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.13)
Proof of Theorem 1.
Using Lemma A.10 with $\mu>0$ and (A.6), we have
$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}].$ (A.14)
Since $\alpha_{k}\equiv\alpha\leqslant 1/(10L(HM+\eta^{2}))\leqslant M/(4L(M+\eta^{2}))$, by Lemma A.8, we have
$\frac{1}{2M}\sum_{m=1}^{M}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{1}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{1}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}+\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
Multiplying the weight $w_{k}$ and summing over $k\in\llbracket 0,K\rrbracket$, and by (A.14), we have
$\quad\,\frac{1}{2M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]$
$\leqslant\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right)+\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right)+\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}],$
which is equivalent to
$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{10}{W_{K}}\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right).$ (A.15)
We observe that $1-\mu\alpha/2\geqslant 1-\mu/(20L(HM+\eta^{2}))\geqslant 1-1/(8HM)$, we choose $w_{k}=(1-\mu\alpha/2)^{-k-1}$ which are $2H$-slow increasing. Then by Lemma A.6, we obtain the desired result. ∎
Proof of Theorem 2.
For the convex case with $\mu=0$, we simply use Lemma A.6 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$. ∎
Proof of Theorem 3.
By Lemma A.9 and (A.6), and define $u_{k}\coloneqq 4\cdot\mathbb{E}[F(\overline{x}_{k})-F^{\star}]$, we have
$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ $\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right)+\frac{2L\alpha_{k}}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
$\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right)+\frac{2\alpha_{k}}{15LMW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2},$
by (A.13). We simplify this to
$\left(1-\frac{2\alpha_{k}}{15L}\right)\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right).$
Applying Lemma A.7 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$ gives the desired result.
Appendix B Details of Numerical Experiments
All experiments are conducted on a Slurm cluster. For experiments with 4 GPUs, each node has either 4 NVIDIA GEFORCE RTX 2080 Ti (11GB) GPUs or 4 NVIDIA A40 (48GB) GPUs or 4 NVIDIA L40s (48GB) GPUs. For experiments with 8 GPUs, each node has 8 NVIDIA A100 (80GB) GPUs. For CPUs, each node has either 64 cores or 96 cores. The total RAM is 512GB or 1TB.
B.1 Three-layer Convolutional Neural Network on MNIST
Table 4: Training hyperparameters for three-layer CNN on MNIST
Model Three-layer CNN on MNIST
Training samples 6M (100 epochs)
Weight initialization Default
Optimizer Adam
Learning rate schedule None
$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
$\varepsilon$ $10^{-8}$
Base learning rate 0.0008
Base batch size 8
Maximum batch size 50000
Weight decay 0
Base global batch size 4096
Base local batch size 1024
Maximum global batch size 60000
Maximum local batch size 15000
Weight decay 0
Figure 6: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies with three different seeds.
Figure 7: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
Figure 8: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
Figure 9: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
B.2 ResNet-50 on CIFAR-10
Table 5: Training hyperparameters for ResNet-50 on CIFAR-10
Model ResNet-50 on CIFAR-10
Training samples 10M (200 epochs)
Weight initialization Default
Optimizer Momentum SGD (SHB)
Learning rate schedule Linear warmup + cosine decay
Learning rate warmup (samples) 1M (10%)
Peak learning rate 0.05
Minimum learning rate 0.005
Base learning rate 0.05
Base global batch size 256
Base local batch size 64
Maximum global batch size 50000
Maximum local batch size 12500
Weight decay 0.0005
Momentum 0.9
Figure 10: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with three different seeds.
Figure 11: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
Figure 12: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
Figure 13: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
Figure 14: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with 8 workers with three different seeds.
Figure 15: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with 8 workers with two other different seeds.
Figure 16: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with 8 workers with three different seeds.
Figure 17: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with 8 workers with three different seeds.

### v2

@@
 Modern deep neural networks often require distributed training with many workers due to their large size.
-As worker numbers increase, communication overheads become the main bottleneck in data-parallel minibatch stochastic gradient methods with per-iteration gradient synchronization.
-Local gradient methods like Local SGD reduce communication by only syncing after several local steps.
-Despite understanding their convergence in i.i.d.
-and heterogeneous settings and knowing the importance of batch sizes for efficiency and generalization, optimal local batch sizes are difficult to determine.
+As the number of workers increases, communication overheads become the main bottleneck in data-parallel minibatch stochastic gradient methods with per-iteration gradient synchronization.
+Local gradient methods like Local SGD reduce communication by only synchronizing model parameters and/or gradients after several local steps.
+Despite an understanding of their convergence and the importance of batch sizes for training efficiency and generalization, optimal batch sizes for local gradient methods are difficult to determine.
 We introduce adaptive batch size strategies for local gradient methods that increase batch sizes adaptively to reduce minibatch gradient variance.
-We provide convergence guarantees under homogeneous data conditions and support our claims with image classification experiments, demonstrating the effectiveness of our strategies in training and generalization.
+We provide convergence guarantees under homogeneous data conditions and support our claims with image classification and language modeling experiments, demonstrating the effectiveness of our strategies for both training efficiency and generalization.
 1 Introduction
-Recent advances in deep learning, particularly large language models (LLMs) with billions of parameters, necessitate efficient distributed training on many GPUs or TPUs.
-Data-parallel optimization [83] simplifies this by having each worker compute and average batch gradients from different i.i.d.
+Recent advances in deep learning, particularly large language models (LLMs) with more than billions of parameters, necessitate efficient distributed training on many workers such as GPUs or TPUs.
+Data-parallel optimization (Zinkevich et al., 2010) simplifies this by having each worker compute and average batch gradients from different i.i.d.
 data batches, followed by a global model update.
 Despite its efficiency with numerous workers, this method faces challenges such as costly gradient synchronization and reduced generalization from large batches.
-Frequent synchronization of models or gradients among workers leads to significant communication overheads, exacerbated as worker counts and model sizes increase [71, 48, 79].
+Frequent synchronization of model parameters or gradients among workers leads to significant communication overheads, exacerbated as worker counts and model sizes increase (Tang et al., 2021; Li et al., 2022; Xu et al., 2023).
 This challenge intensifies with workers distributed across nodes with limited inter-node communication speeds.
-To mitigate this, communication-efficient distributed optimization methods [43], such as local gradient methods [67, 82, 81, 26, 73] and their variants, have been developed.
-In local stochastic gradient methods (Local SGD), workers train locally and synchronize less frequently by averaging models after several local steps, effectively reducing communication overheads.
+To mitigate this, communication-efficient distributed optimization methods (Lan et al., 2020), such as local gradient methods (Stich, 2019a; Zhou and Cong, 2018; Yu et al., 2019; Haddadpour et al., 2019; Wang and Joshi, 2019) and their variants, have been developed.
+In local stochastic gradient methods (Local SGD), workers train locally and synchronize less frequently by averaging model parameters after several local steps, effectively reducing communication overheads.
 In data-parallel optimization, the total batch size is the sum of all workers’ local batch sizes, increasing with more workers.
-Large batch sizes are preferred to fully utilize worker memory, making large-batch training the standard in large-scale model training.
-However, this approach often results in reduced model generalization, a problem noted by LeCun et al.
-[47].
-Large-batch training typically creates a loss landscape with sharp minima, worsening generalization [35].
-Although increasing the learning rate proportionally to batch size can mitigate the generalization gap [29, 65, 66], it cannot fully close it.
+Attributed to the ever-rising memory of workers, large batch sizes are preferred to fully utilize worker memory, making large-batch training the de facto paradigm in large-scale model training.
+However, this approach often results in worse model generalization, a problem noted by LeCun et al.
+(2002).
+Large-batch training typically creates a loss landscape with sharp minima, worsening generalization (Keskar et al., 2017).
+Although scaling the learning rate with the batch size can mitigate the so-called generalization gap (Hoffer et al., 2017; Smith and Le, 2018; Smith et al., 2018), it cannot fully close it.
 Keskar et al.
-[35] recommended adaptive sampling to gradually increase batch sizes during training [7, 19], explored in [11, 12, 44, 59].
+(2017) recommended adaptive sampling to gradually increase batch sizes during training (Byrd et al., 2012; Friedlander and Schmidt, 2012), explored in De et al.
+(2016, 2017); Lau et al.
+(2024); Ostroukhov et al.
+(2024).
 However, this method is still confined to single- or few-worker scenarios and does not scale well with larger models and more workers.
 Motivated by the above two challenges, we answer the following question affirmatively:
-Can we develop training strategies which are communication-efficient and memory-efficient with desirable generalization performance for large-scale distributed training?
+Can we develop communication-efficient and memory-efficient training strategies with desirable generalization performance for large-scale distributed training?
 Contributions.
-We make the following contributions: (i) We develop adaptive batch size strategies for local gradient methods by extending the adaptive sampling methods in [7, 3].
+We make the following contributions: (i) We develop adaptive batch size strategies for local gradient methods by extending the adaptive sampling methods in Byrd et al.
+(2012); Bollapragada et al.
+(2018).
 Such strategies are for local batch sizes, i.e., batch sizes of individual workers.
@@
 (ii) Under various standard assumptions, we provide convergence guarantees of the proposed approaches for smooth (strongly) convex and nonconvex objectives under a homogeneous data setting.
-(iii) We empirically demonstrate the efficacy of the proposed strategies with a three-layer convolutional neural network and ResNet for image classification on the MNIST and CIFAR-10 datasets, respectively, in terms of communication efficiency, memory efficiency, and model generalization performance.
+(iii) We empirically demonstrate the efficacy of the proposed strategies with ResNet for image classification on the CIFAR-10 and ImageNet datasets and MicroLlama on the C4 dataset respectively, in terms of communication efficiency, memory efficiency, and model generalization performance.
 To grasp a thorough understanding of the strategies, we also study the effect of different hyperparameters such as the number of local steps and the one which controls the probability of increasing batch sizes.
@@
 Minibatch gradient methods.
-Minibatch SGD [13, 63, 4, 15] is the simplest and most widely used optimizer for large-scale machine learning [5], and is easily extendable with data parallelism [49].
-The convergence of minibatch SGD and its variants like AdaGrad [18] and Adam [39] is well-understood across various settings, including convex [21], nonconvex [36], and nonsmooth [10] objectives, under different gradient variance assumptions.
+Minibatch SGD (Dean et al., 2012; Robbins and Monro, 1951; Bottou, 2010; Dekel et al., 2012) is the simplest and most widely used optimizer for large-scale machine learning (Bottou et al., 2018), and is easily extendable with data parallelism (Li et al., 2020).
+The convergence of minibatch SGD and its variants like AdaGrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2015) is well-understood across various settings, including convex (Gower et al., 2019), nonconvex (Khaled and Richtárik, 2023), and nonsmooth (Davis et al., 2020) objectives, under different gradient variance assumptions.
 However, frequent gradient synchronization among workers leads to communication overheads, posing a significant challenge in distributed training.
 Local gradient methods.
-Local SGD [67, 83, 54], a communication-efficient variant of minibatch SGD, reduces overheads from frequent synchronization in distributed training.
-Recent studies [51, 17] highlight its efficacy in pre-training large language models with extensive computational resources.
-Optimization of Local SGD for varied data settings has shown promising convergence rates [67, 77, 76, 40, 37, 69, 74].
-It includes extensions like asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75].
-Federated learning [54, 9, 33, 34, 62], another model focusing on efficiency, addresses heterogeneity and partial participation.
+Local SGD (Stich, 2019a; Zinkevich et al., 2010; McMahan et al., 2017), a communication-efficient variant of minibatch SGD, reduces overheads from frequent synchronization in distributed training.
+Recent studies (Liu et al., 2024; Douillard et al., 2023) highlight its efficacy in pre-training large language models with extensive computational resources.
+Optimization of Local SGD for varied data settings has shown promising convergence rates (Stich, 2019a; Woodworth et al., 2020b, a; Koloskova et al., 2020; Khaled et al., 2020; Stich and Karimireddy, 2020; Wang and Joshi, 2021).
+It includes extensions like asynchronous SGD (Mishchenko et al., 2022; Koloskova et al., 2022; Nguyen et al., 2022; Leconte et al., 2024; Islamov et al., 2024) and hierarchical SGD (Lin et al., 2020; Castiglia et al., 2021; Wang et al., 2022).
+Federated learning (McMahan et al., 2017; Chen et al., 2021; Kairouz et al., 2021; Karimireddy et al., 2020; Reddi et al., 2021), another model focusing on efficiency, addresses heterogeneity and partial participation.
 While applicable, we defer exploring federated learning’s theoretical and empirical aspects to future work.
@@
 Large-batch training not only faces communication overheads in minibatch SGD under distributed data parallelism but also shows poorer generalization despite better hardware efficiency.
-It has been successful in tasks like ImageNet classification and BERT training [80, 22, 64].
+It has been successful in tasks like ImageNet classification and language model training (You et al., 2020; Goyal et al., 2017; Akiba et al., 2017; Shallue et al., 2019; Liu et al., 2019; Puri et al., 2018).
 Addressing these, Lin et al.
-[50] introduced Post-local SGD, a hybrid of minibatch and Local SGD, where early training involves frequent synchronization of local models, followed by Local SGD.
+(2020) introduced Post-local SGD, a hybrid of minibatch and Local SGD, where early training involves frequent synchronization of local models, followed by Local SGD.
 In large-scale environments, Ortiz et al.
-[58] noted that Post-local SGD does not consistently outperform Local SGD in generalization, and the transition from minibatch to Local SGD presents a trade-off between communication efficiency and generalization.
+(2021) noted that Post-local SGD does not consistently outperform Local SGD in generalization, and the transition from minibatch to Local SGD presents a trade-off between communication efficiency and generalization.
 Following these findings, Gu et al.
-[25, 24] analyzed Local SGD’s generalization via an SDE approximation and proposed the Quadratic Synchronization Rule (QSR) for adjusting local gradient steps during training.
+(2024, 2023) analyzed Local SGD’s generalization via an SDE approximation and proposed the Quadratic Synchronization Rule (QSR) for adjusting local gradient steps during training.
 Adaptive batch size strategies.
-In stochastic optimization, adaptive sampling methods [7, 19, 3] adjust batch sizes based on gradient noise, enhancing training processes for smooth finite-sum problems.
-These methods, further explored in deep learning [11, 12, 44, 59], have not been applied to data parallelism with distributed batches.
-The strategies discussed here draw from batch size scheduling, a key technique in pretraining LLMs, evidenced by its use in GPT-3 [6], Nemotron-4 [60], OLMo 65B [23], and DeepSeek-V2 [14], which involve increasing batch sizes in set stages.
+In stochastic optimization, adaptive sampling methods (Byrd et al., 2012; Friedlander and Schmidt, 2012; Bollapragada et al., 2018) adjust batch sizes based on gradient noise, enhancing training processes for smooth finite-sum problems.
+These methods, further explored in deep learning (De et al., 2016, 2017; Lau et al., 2024; Ostroukhov et al., 2024), have not been applied to data parallelism with distributed batches.
+The strategies discussed here draw from batch size scheduling, a key technique in pretraining LLMs, evidenced by its use in GPT-3 (Brown et al., 2020), Nemotron-4 (Parmar et al., 2024), OLMo 65B (Groeneveld et al., 2024), and DeepSeek-V2 (DeepSeek-AI, 2024), which involve increasing batch sizes in set stages.
 Despite their benefits for efficiency and parallelization, these strategies are heuristic and have unclear impacts on training.
-This work aims to develop theoretically grounded, communication-efficient adaptive batch size strategies akin to adaptive sampling, viewed as stochastic variance-reduction methods [31, 78, 20], focusing on scalable batch size increments.
-Several studies have explored adaptive batch size strategies in deep learning, including Big Batch SGD [11, 12], CABS [1], AdaBatch [16], and SimiGrad [61].
-Conversely, AdaScale SGD [32] modifies learning rates for large batches instead of batch sizes.
+This work aims to develop theoretically grounded, communication-efficient adaptive batch size strategies akin to adaptive sampling, viewed as stochastic variance-reduction methods (Johnson and Zhang, 2013; Xiao and Zhang, 2014; Gower et al., 2020), focusing on scalable batch size increments.
+Several studies have explored adaptive batch size strategies in deep learning, including Big Batch SGD (De et al., 2016, 2017), CABS (Balles et al., 2017), AdaBatch (Devarakonda et al., 2017), and SimiGrad (Qin et al., 2021).
+Conversely, AdaScale SGD (Johnson et al., 2020) modifies learning rates for large batches instead of batch sizes.
 These methods, however, are often not theoretically robust, lacking solid convergence proofs or are restricted to SGD analyses under strict conditions like convexity or the Polyak–Łojasiewicz condition.
-Their batch size adjustment strategies usually depend on heuristic rules, such as geometric growth or decay [61], which can lead to unreliable outcomes.
-Crucially, these methods overlook the distributed data-parallel context, prevalent in practice, thus not addressing the needs of adaptive batch size strategies in large-scale distributed training.
+Their batch size adjustment strategies usually depend on heuristic rules, such as geometric growth or decay (Qin et al., 2021), which can lead to unreliable outcomes due to the lack of theoretical guarantees.
+Crucially, these methods overlook the distributed data-parallel context, prevalent in practice, thus failing to address the needs of adaptive batch size strategies in large-scale distributed training.
 Hyperparameter tuning for Local SGD.
-Besides batch sizes, recent research has also focused on optimizing learning rate schedules for Local SGD, particularly in identical data settings as discussed in [2].
-This optimization is crucial in federated learning with heterogeneous data, where local objectives differ significantly in geometry and smoothness, necessitating locally adaptive learning rate schedules [56, 38].
+Besides batch sizes, recent research has also focused on optimizing learning rate schedules for Local SGD, particularly in identical data settings as discussed in Balles et al.
+(2024).
+This optimization is crucial in federated learning with heterogeneous data, where local objectives differ significantly in geometry and smoothness, necessitating locally adaptive learning rate schedules (Mukherjee et al., 2024; Kim et al., 2024).
 Our adaptive local batch size strategies complement these efforts.
@@
 We consider a distributed training setting with $M$ workers, each with a possibly heterogeneous (i.e.
-non-i.i.d.
-) underlying true data distribution $\mathbb{P}_{m}$ ($m\in\llbracket M\rrbracket$).
+non-i.i.d.) underlying true data distribution $\mathbb{P}_{m}$ ($m\in\llbracket M\rrbracket$).
 The true data distribution $\mathbb{P}_{m}$ is usually unknown, but can be approximated by its empirical distribution ^$\widehat{\mathbb{P}}_{m}=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}\updelta_{\xi_{i}^{m}}$, where $\mathcal{D}_{m}\coloneqq\{\xi_{i}^{m}\}_{i\in\llbracket n_{m}\rrbracket}$ is the set of $n_{m}$ data samples of worker $m$ and $\xi_{i}^{m}\in\mathcal{Z}_{m}\subseteq\mathbb{R}^{p}$ for $i\in\llbracket n_{m}\rrbracket$.
@@
 This is followed by the global model update, performed via $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$, where $\alpha_{k}$ is the learning rate of iteration $k$.
-The same mechanism also applies to other optimizers such as Adam [39] and AdaGrad [18].
+The same mechanism also applies to other optimizers such as Adam (Kingma and Ba, 2015) and AdaGrad (Duchi et al., 2011).
 Local SGD.
-As opposed to minibatch SGD, Local SGD [67] reduces the communication frequency.
-Communication for model averaging (through all-reduce operations) is performed every $H$ local gradient steps, Local SGD takes the following updates:
-$(\forall m\in\llbracket M\rrbracket)(\forall k\in\mathbb{N})\quad x_{k+1}^{m}=\begin{cases}\frac{1}{M}\sum_{m=1}^{M}\left(x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right)&\text{if }H\text{ divides }k+1,\\ x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})&\text{otherwise}.\end{cases}$ (2)
+As opposed to minibatch SGD, Local SGD (Stich, 2019a) reduces the communication frequency.
+Communication for model averaging (through all-reduce operations) is performed every $H$ local gradient steps, Local SGD takes the following updates: for $m\in\llbracket M\rrbracket)$ and $k\in\mathbb{N}$,
+$x_{k+1}^{m}=\begin{cases}\frac{1}{M}\sum_{m=1}^{M}\left(x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right)&\text{if }H|k+1,\\ x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})&\text{otherwise}.\end{cases}$ (2)
 We also consider the following alternative representation that explicitly separates the numbers of local gradient steps $H$ and communication rounds $K$.
-In particular, for $M$ parallel workers, $K$ rounds of communication, $H$ local gradient steps per round, Local SGD updates can also be expressed as
-$(\forall m\in\llbracket M\rrbracket)(\forall k\in\llbracket 0,K-1\rrbracket)\;\begin{cases}x_{k,h+1}^{m}=x_{k,h}^{m}-\alpha_{k,h}\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})&\text{for }h\in\llbracket 0,H-1\rrbracket,\\ x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}.\end{cases}$ (3)
+In particular, for $M$ parallel workers, $K$ rounds of communication, $H$ local gradient steps per round, Local SGD updates can also be expressed as: for $m\in\llbracket M\rrbracket)$ and $k\in\llbracket 0,K-1\rrbracket$,
+$\begin{cases}x_{k,h+1}^{m}=x_{k,h}^{m}-\alpha_{k,h}\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})&\text{for }h\in\llbracket 0,H-1\rrbracket,\\ x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}.\end{cases}$ (3)
 4 Adaptive Local Batch Size Strategies for Local Gradient Methods
-The proposed adaptive local batch size strategies are based on the adaptive sampling method [7] originally developed for unconstrained stochastic optimization problems using minibatch SGD.
+The proposed adaptive local batch size strategies are based on the adaptive sampling method (Byrd et al., 2012) originally developed for unconstrained stochastic optimization problems using minibatch SGD.
 We outline its mechanism and its extension to Local SGD below.
 4.1 Adaptive Sampling Methods
-Introduced in [7], the adaptive sampling method for minibatch SGD was designed for a single worker ($M=1$) with the objective $F(x)=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i})$ over samples $\mathcal{D}=\{\xi_{i}\}_{i=1}^{n}$.
+Introduced in Byrd et al.
+(2012), the adaptive sampling method for minibatch SGD was designed for a single worker ($M=1$) with the objective $F(x)=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i})$ over samples $\mathcal{D}=\{\xi_{i}\}_{i=1}^{n}$.
 This method relies on a key stochastic optimization principle: when the sample size $n$ is large, the batch gradient $\nabla F_{\mathcal{B}_{k}}$ approximates the full gradient $\nabla F$.
@@
 $\frac{1}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))\leqslant\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}.$ (6)
-In practice, for each iteration $k\in\mathbb{N}$, the dynamic sample gradient algorithm [7] performs the following minibatch SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with learning rate $\alpha_{k}>0$, and then checks the following (approximate) norm test condition (6), which can also be viewed as an approximation of the exact variance norm test in the stochastic setting:
+In practice, for each iteration $k\in\mathbb{N}$, the dynamic sample gradient algorithm (Byrd et al., 2012) performs the following minibatch SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with learning rate $\alpha_{k}>0$, and then checks the following (approximate) norm test condition (6), which can also be viewed as an approximation of the exact variance norm test in the stochastic setting:
 $\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|^{2}\right]\leqslant\eta^{2}\|\nabla F(x_{k})\|^{2},$ (7)
@@
 Here we abbreviate the conditional expectation on $\mathcal{F}_{k}$ (i.e., the filtration $\mathcal{F}_{k}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$.
-If condition (6) is violated, the batch size of the next iteration $b_{k+1}$ is increased, determined by the formula:
-$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\rceil,$
+If condition (6) is violated, the batch size of the next iteration $b_{k+1}$ is increased, determined by:
+$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\rceil,$ (8)
 and the next batch $\mathcal{B}_{k+1}$ is sampled with this larger batch size accordingly.
@@
 To moderate these schedules, Bollapragada et al.
-[3] introduced the (augmented) inner product test, which manages the variance between the batch and full gradient’s inner product.
+(2018) introduced the (augmented) inner product test, which manages the variance between the batch and full gradient’s inner product.
 Given the preference for large batch sizes for memory efficiency, controlled via $\ eta$, we focus on the norm test and defer exploring local variants of the (augmented) inner product test to future work.
@@
 We extend the above adaptive sampling methods beyond the single worker setting with data parallelism using multiple workers, more specifically for Local SGD.
-We have to invoke a local variant of the norm test (7), called the (exact variance) local norm test, at every iteration $k\in\mathbb{N}$, using formulation (2): for some sequence of positive constants $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$,
-$(\forall m\in\llbracket M\rrbracket)\quad\mathbb{E}_{\mathcal{F}_{k}^{m}}\left[\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})-\nabla F_{m}(x_{k}^{m})\|^{2}\right]\leqslant\eta_{m}^{2}\left\lVert\nabla F_{m}(x_{k}^{m})\right\rVert^{2},$ (8)
+We have to invoke a local variant of the norm test (7), called the (exact variance) local norm test, at every iteration $k\in\mathbb{N}$, using formulation (2): for some sequence of positive constants $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$, for all $m\in\llbracket M\rrbracket$,
+$\mathbb{E}_{\mathcal{F}_{k}^{m}}\left[\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})-\nabla F_{m}(x_{k}^{m})\|^{2}\right]\leqslant\eta_{m}^{2}\left\lVert\nabla F_{m}(x_{k}^{m})\right\rVert^{2},$ (9)
 where $\mathbb{E}_{\mathcal{F}_{k}^{m}}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}^{m}]$ denotes the conditional expectation on $\mathcal{F}_{k}^{m}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0}^{m},\ldots,\mathcal{B}_{k-1}^{m}\})$.
 Again, $\nabla F_{m}$ is in general unavailable since the sample size $n_{m}$ could be large.
-The above local norm test is approximately enforced by the following (approximate) local norm test:
-$(\forall m\in\llbracket M\rrbracket)\quad\frac{1}{b_{k}^{m}}\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}\left(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m})\right)\leqslant\eta_{m}^{2}\left\lVert\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right\rVert^{2}.$ (9)
+The above local norm test is approximately enforced by the following (approximate) local norm test: for all $m\in\llbracket M\rrbracket$,
+$\frac{1}{b_{k}^{m}}\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}\left(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m})\right)\leqslant\eta_{m}^{2}\left\lVert\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right\rVert^{2}.$ (10)
 Should the condition (10) not be met, the subsequent local batch sizes are determined by the formula:
@@
 and the local batches $\mathcal{B}_{k+1}^{m}$ are augmented accordingly.
-In a practical implementation, for simplicity, we can set $\eta_{m}\equiv\eta$ for all $m\in\llbracket M\rrbracket$ and choose the next local batch sizes to be
-$(\forall m\in\llbracket M\rrbracket)\quad b_{k+1}^{m}\equiv b_{k+1}\coloneqq\max_{m\in\llbracket M\rrbracket}\ \left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\|^{2}}\right\rceil,$ (11)
+In a practical implementation, for simplicity, we can set $\eta_{m}\equiv\eta$ for all $m\in\llbracket M\rrbracket$ and choose the next local batch sizes to be, for all $m\in\llbracket M\rrbracket$,
+$b_{k+1}^{m}\coloneqq\max_{m\in\llbracket M\rrbracket}\ \left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}(\nabla f_{m}(x_{k}^{m};\xi_{i}^{m}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\|^{2}}\right\rceil,$
 which is particularly useful when homogeneous workers and identical data on the workers are used, so that there would not be discrepancy of training time for the local steps due to different local batch sizes and undesirable stragglers before model averaging.
-The proposed adaptive batch size strategies extend to local variants of minibatch stochastic gradient optimizers beyond SGD [63], including momentum SGD [70], AdaGrad [18], Adam [39], and AdamW [52].
-The algorithm is detailed in Algorithm 1.
-In a data parallel setting, adaptive global batch size strategies can be developed using minibatch SGD, leading to increased communication overheads from aggregating local batch gradient variances among workers to compute the quantities in (6).
-Algorithm 1 Adaptive Local Batch Size Schedules for Local Gradient Methods
-no.
-of workers $M\in\mathbb{N}^{*}$, no.
-of communication steps $K\in\mathbb{N}^{*}$, no.
-of local steps $H\in\mathbb{N}^{*}$, no.
-of training samples $N\in\mathbb{N}^{*}$, global step counter $k=0$, processed samples counter $B=0$, training sets $(\mathcal{D}_{m})_{m\in\llbracket M\rrbracket}=(\{\xi_{i}^{m}\}_{i=1}^{n_{m}})_{m\in\llbracket M\rrbracket}$, initial model parameters $x_{0,0}^{m}=x_{0}\in\mathbb{R}^{d}$for all $m\in\llbracket M\rrbracket$, initial local batch sizes $(b_{0}^{m})_{m\in\llbracket M\rrbracket}$, learning rate schedules $(\alpha_{k,h})_{k\in\llbracket 0,K-1\rrbracket,h\in\llbracket 0,H-1\rrbracket}$, $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$$B<N$worker $m=1,\ldots,M$in parallel local step $h=0,\ldots,H-1$
-Sample the local batch $\mathcal{B}_{k,h}^{m}$ of size $b_{k,h}^{m}\coloneqq|\mathcal{B}_{k,h}^{m}|$ uniformly from $\mathcal{D}_{m}$
-Compute per-sample gradients $\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$ for each $i\in\mathcal{B}_{k,h}^{m}$
-Compute the local batch gradient $\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})=\frac{1}{b_{k,h}^{m}}\sum_{i\in\mathcal{B}_{k,h}^{m}}\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$
-Compute $\mathrm{Var}_{i\in\mathcal{B}_{k,h}^{m}}(\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m}))$ by applying (5)
-Compute the local norm test statistic $\mathsf{T}_{k,h}^{m}$ defined by (10)
-Determine the next local batch sizes: $b_{k,h+1}^{m}=\max\{\mathsf{T}_{k,h}^{m},b_{k,h}^{m}\}$
-[…]
-Update the global model: $x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}$ for each $m\in\llbracket M\rrbracket$triangleright all-reduce
-$B\leftarrow B+\sum_{m=1}^{M}\sum_{h=0}^{H-1}b_{k,h}^{m}$
-[…]
-[…]
-[…]
-[…]
-[…]
+The proposed adaptive batch size strategies extend to local variants of minibatch stochastic gradient optimizers beyond SGD (Robbins and Monro, 1951), including momentum SGD (Sutskever et al., 2013), AdaGrad (Duchi et al., 2011), Adam (Kingma and Ba, 2015), and AdamW (Loshchilov and Hutter, 2019).
+In a data-parallel setting, adaptive global batch size strategies can be developed using minibatch SGD, leading to increased communication overheads from aggregating local batch gradient variances among workers to compute the quantities in (6).
+4.3 Practical Considerations
+The (approximate) norm test (6) and its local variant (10) extend the norm test (Byrd et al., 2012) for Local SGD, but implementing them in deep learning libraries like PyTorch (Paszke et al., 2019) is complex, as only batch gradients $\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$, not per-sample gradients $\nabla f_{m}(x_{k}^{m};\xi_{i}^{m})$, are available in the backward training step.
+Therefore, we cannot compute the variance of per-sample gradients without additional subroutines (e.g., using torch.func).
+These methods do not scale well to large models, as they require a full model replica and per-sample gradients matching the model’s parameter size.
+Increasing batch sizes during training elevates memory demands, necessitating many gradient accumulation steps (a serial computation) to simulate large batches.
+We, therefore, suggest an alternative to approximate the variance term in the (local) norm test.
+We can use local batch gradients as estimators instead of per-sample gradients.
+To illustrate, consider minibatch SGD with data parallelism among multiple workers where $\nabla F_{\mathcal{B}_{k}}(x_{k})=\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$ and $x_{k}^{m}=x_{k}$ apply for $m\in\llbracket M\rrbracket$.
+Note that $\mathrm{Var}_{m\in\llbracket M\rrbracket}\left(\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})\right)\coloneqq\frac{1}{M-1}\sum_{m=1}^{M}\left\lVert\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})-\nabla F_{\mathcal{B}_{k}}(x_{k})\right\rVert^{2}$.
+Then, by the definition of $\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$, we can write
+$\mathrm{Var}_{m\in\llbracket M\rrbracket}\left(\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})\right)=\sum_{i\in\mathcal{B}_{k}^{m}}\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}^{m}}\left(\nabla f(x_{k};\xi_{i}^{m})\right)}{(b_{k}^{m})^{2}}=\frac{1}{b_{k}^{m}}\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k};\xi_{i})\right)=\frac{M}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k};\xi_{i})\right),$
+since $b_{k}^{m}=b_{k}/M$.
+Therefore, for the norm test (8), we can instead compute $\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k};\xi_{i})\right)=b_{k}/M\cdot\frac{1}{M-1}\sum_{m=1}^{M}\left\lVert\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})-\nabla F_{\mathcal{B}_{k}}(x_{k})\right\rVert^{2}$.
+We extend this workaround to Local SGD, focusing on the i.i.d.
+setting where ^$\widehat{\mathbb{P}}_{m}\equiv\widehat{\mathbb{P}}$ and $f_{m}\equiv f$, thus resulting in $F_{m}=F$ for all $m\in\llbracket M\rrbracket$.
+Varying $x_{k}^{m}$’s across workers $m\in\llbracket M\rrbracket$ due to different $\mathcal{B}_{k}^{m}$’s prevent local gradients from being evaluated at the same iterates as in minibatch SGD during updates.
+Thus, we instead compute
+$\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k}^{m};\xi_{i}^{m})\right)=\frac{b_{k}}{M}\cdot\frac{1}{M-1}\sum_{m=1}^{M}\left\lVert\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})-\overline{g}_{k}\right\rVert^{2},$
+where $\overline{g}_{k}\coloneqq\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$.
+The actual (exact variance) norm test for local gradient methods is given by, for all $m\in\llbracket M\rrbracket$,
+$\mathbb{E}_{\mathcal{F}_{k}^{m}}\left[\|\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})-\nabla F(x_{k}^{m})\|^{2}\right]\leqslant\eta_{m}^{2}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2},$ (12)
+In our implementation, the above actual (exact variance) norm test for local gradient methods is approximately enforced by the following (approximate) norm test for local gradient methods:
+$\frac{1}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k}^{m};\xi_{i}^{m})\right)\leqslant\eta^{2}\left\lVert\overline{g}_{k}\right\rVert^{2}.$ (13)
+Should the condition (13) not be met, the subsequent local batch sizes are determined by the formula:
+$(\forall m\in\llbracket M\rrbracket)\;b_{k+1}^{m}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k}^{m};\xi_{i}^{m})\right)}{M\eta^{2}\left\lVert\overline{g}_{k}\right\rVert^{2}}\right\rceil,$ (14)
+and the local batches $\mathcal{B}_{k+1}^{m}$ are augmented accordingly.
+Note that the computation of $\overline{g}_{k}$ requires additional all-reduce communication of $g_{k}^{m}$ among the $M$ workers.
+As a result, to avoid unnecessary additional communication rounds due to the test, we only perform the test every $H$ local gradients steps, i.e., performing model parameter and gradient synchronization at the same time and frequency.
+The algorithm pseudocode is detailed in Appendix A.
 5 Convergence Analysis
 Our analysis focuses on the i.i.d.
-setting as shown in (1).
-We set $\eta_{m}\equiv\eta\in(0,1)$ for all $m\in\llbracket M\rrbracket$, typical in datacenter environments for distributed training [13], where workers access the entire data set.
+setting where ^$\widehat{\mathbb{P}}_{m}\equiv\widehat{\mathbb{P}}$ and $f_{m}\equiv f$, and hence $F_{m}=F$ for all $m\in\llbracket M\rrbracket$.
+We set $\eta_{m}\equiv\eta\in(0,1)$ for all $m\in\llbracket M\rrbracket$, typical in datacenter environments for distributed training (Dean et al., 2012), where workers access the entire data set.
 The analysis of heterogeneous data settings is deferred to future work.
-Our convergence analysis, following [69], does not assume bounded gradient variance across $\mathbb{R}^{d}$.
-Instead, our adaptive local batch size strategies ensure the expected strong growth condition [36, 5] at each local iterate $x_{k}^{m}$ for all workers $m\in\llbracket M\rrbracket$.
-This approach relaxes the standard assumptions of uniform boundedness and expected strong growth conditions [72], which are challenging to verify except in the interpolation regime [53].
+Our convergence analysis, following Stich and Karimireddy (2020), does not assume bounded gradient variance across $\mathbb{R}^{d}$.
+Instead, our adaptive local batch size strategies ensure the expected strong growth condition (Khaled and Richtárik, 2023; Bottou et al., 2018) at each local iterate $x_{k}^{m}$ for all workers $m\in\llbracket M\rrbracket$.
+This approach relaxes the standard assumptions of uniform boundedness and expected strong growth conditions (Vaswani et al., 2019), which are challenging to verify except in the interpolation regime (Ma et al., 2018).
 We might need to invoke the following assumptions on the objective function $F$.
@@
 If $\mu=0$, we say that $F$ is convex.
-We derive the following convergence results for the proposed adaptive batch size strategies as follows. Their proofs are deferred to Appendix A.
+We derive the following convergence results for the proposed adaptive batch size strategies as follows. Their proofs are deferred to Appendix B.
 Theorem 1 (Strongly convex; $\mu>0$).
@@
 Then, for some $\alpha\leqslant 1/(10L(HM+\eta^{2}))$, we have
-$\mathbb{E}[F(x_{\mathsf{out}})]-F^{\star}=\widetilde{\mathscr{O}}\left(L(HM+\eta^{2})\exp\left\{-\frac{\mu K}{10L(HM+\eta^{2})}\right\}\left\lVert x_{0}-x^{\star}\right\rVert^{2}\right),$ (12)
+$\mathbb{E}[F(x_{\mathsf{out}})]-F^{\star}=\widetilde{\mathscr{O}}\left(L(HM+\eta^{2})\mathrm{e}^{-\frac{\mu K}{10L(HM+\eta^{2})}}\left\lVert x_{0}-x^{\star}\right\rVert^{2}\right),$ (15)
 where $x_{\mathsf{out}}\in\{x_{k}^{m}\}_{k\in\llbracket 0,K-1\rrbracket,m\in\llbracket M\rrbracket}$ is chosen as $x_{k}^{m}$ with probability proportional to $(1-\mu\alpha/2)^{-k}$ uniformly on $m\in\llbracket M\rrbracket$ and $\widetilde{\mathscr{O}}$ hides logarithmic factors.
@@
 The theorems confirm that Local SGD, using the local norm test with exact variance, achieves linear convergence rates for both convex and nonconvex objectives.
-This scaling is proportional to the number of local gradient steps $H$ and workers $M$, akin to Local SGD under constant batch sizes and bounded gradient variance [69].
+This scaling is proportional to the number of local gradient steps $H$ and workers $M$, akin to Local SGD under constant batch sizes and bounded gradient variance (Stich and Karimireddy, 2020).
 Remark 1 (Choice of $\eta$).
 Convergence rates depend quadratically on $\eta$, which controls the likelihood and size of increased local batch sizes.
-Choosing smaller $\eta$ values accelerates convergence but causes quick batch size escalation, undermining adaptive batch sizing and widening the generalization gap.
+Choosing smaller $\eta$ values accelerates convergence but causes quick batch size escalation, defeating the original purpose of adaptive batch sizing and widening the generalization gap.
 6 Numerical Experiments
-We evaluate the efficiency of the training (total steps or wall-clock time) and the generalization performance of our strategies through image classification tasks.
-We use a three-layer convolutional neural network on the MNIST dataset [46] and a ResNet-50 [28] on the CIFAR-10 dataset [42], with four workers ($M=4$ GPUs).
-The models are trained using Local Adam and Local SGD with momentum (SHB), synchronizing only the models every $H$ steps.
-Further details and results are in Appendix B.
-6.1 Three-layer CNN on MNIST
-We demonstrate our strategies using a three-layer CNN for MNIST classification.
-The training involves 12 million samples (200 epochs) with global and local batch sizes set to 60,000 and 15,000, respectively.
-This task uses a fixed learning rate of 0.0008 without any rate schedules.
-Results are in Table 1.
-Table 1: Adaptive local batch size strategies for three-layer CNN on MNIST; no.
-of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
-Schedule $H=16$ $H=4$ $H=1$
-steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
-Constant 1456 5.01 2048 99.10 1464 9.44 2048 99.06 1464 25.58 2048 99.08
-Constant 352 6.84 8192 98.80 364 8.19 8192 98.85 366 12.37 8192 98.94
-Constant 192 10.59 15000 98.40 196 11.41 15000 98.43 199 13.88 15000 98.54
-$\eta=0.25$ 213 11.35 13414 98.51 221 11.94 13427 98.55 223 14.50 13467 98.66
-$\eta=0.375$ 224 11.85 13077 98.61 228 12.16 13031 98.58 228 14.86 13078 98.71
-$\eta=0.50$ 229 11.85 12615 98.58 235 11.73 12648 98.60 235 14.42 12748 98.72
-Figure 1: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies ($H=16$).
+We evaluate the efficiency of the training (total steps or wall-clock time) and the generalization performance of our strategies through two standard deep learning applications in two domains—image classification and language modeling tasks.
+We use a ResNet-50 (He et al., 2016) on the CIFAR-10 dataset (Krizhevsky, 2009) and a ResNet-101 on the ImageNet dataset (Russakovsky et al., 2015, see Section C.3) for image classification, and a MicroLlama 300M (Wang, 2024) on the C4 dataset (Raffel et al., 2020) for language modeling, with four workers ($M=4$ GPUs).
+The models are trained using Local AdamW and Local SGD with momentum (a.k.a.
+stochastic heavy ball; SHB), synchronizing only the models every $H$ local steps.
+Further experimental details and results are given in Appendix C.
+When the chosen batch sizes cannot fit in the GPU memory, we use the technique of gradient accumulation.
+6.1 ResNet-50 on CIFAR-10
+We fit ResNet-50 for CIFAR-10 using $M=4$ workers, training over 30 million samples (600 epochs) with global and local batch sizes of 50,000 and 12,500.
+Momentum SGD (Sutskever et al., 2013, a.k.a.
+stochastic heavy ball) serves as the inner optimizer.
+Learning rates are managed with a 10% linear warmup and cosine decay, peaking at 0.05 and bottoming out at 0.005.
+Additionally, we apply linear scaling of learning rates (Krizhevsky, 2014; Goyal et al., 2017) relative to batch sizes for constant batch size baselines.
+(a) $H=32$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(b) $H=16$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(c) $H=4$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(d) Adaptive batch size strategies for $H\in\{32,16,4\}$
+Figure 1: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies for ResNet-50 on CIFAR-10.
 Values of $\eta$.
-Table 1 shows that adaptive local batch size strategies can match or exceed the validation accuracy of constant batch size strategies.
-Specifically, with 16 local gradient steps and $\eta=0.25$, the adaptive strategy achieves an average batch size similar to a constant batch size of 15,000 (see Figure 1), but with superior accuracy and 21 additional steps.
-This confirms that adaptive strategies can mitigate the generalization gap in large-batch training.
-Furthermore, Figure 2 indicates that these strategies reach a maximum local batch size of 15,000 later in the training, enhancing hardware utilization and efficiency.
-Similar trends are noted for $H=4$ and $H=1$.
-Figure 2: Validation accuracy and local batch sizes of Local Adam with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
+Table 1 demonstrates that adaptive local batch sizes offer better validation accuracy than constant batch sizes with similar averaged batch size.
+For example, using $H=16$ and $\eta=0.8$, the adaptive approach averages a batch size of 8906, more than double a constant size of 4096, achieving over 1% higher accuracy and nearly 1000 fewer steps.
+This highlights the effectiveness of adaptive strategies in reducing the generalization gap in large-batch training, though it requires 16% more training time due to extra computations from the norm test.
+Similar results are observed with other local gradient step values $H$.
 Local gradient steps $H$.
-Smaller $H$ increases training time due to more frequent communication, and contrary to expectations, does not always enhance generalization.
-As Table 1 demonstrates, with a local batch size of 2048, $H=1$ and $H=4$ have lower accuracy compared to $H=16$.
-This pattern is consistent with adaptive batch sizes, where $H=4$ underperforms $H=16$ at $\eta=0.375$.
-Moreover, while $H=1$ leads to longer training times, it achieves higher accuracy.
-6.2 ResNet-50 on CIFAR-10
-We scale our experiment to ResNet-50 for CIFAR-10, using a training budget of 12 million samples (200 epochs) and setting maximum global and local batch sizes of 60,000 and 15,000, respectively.
-Our learning rate schedules include a 10% linear warmup and cosine decay, with peak and minimum rates of 0.5 and 0.05, respectively.
-Table 2: Adaptive local batch size strategies for ResNet-50 on CIFAR-10; no.
-of total steps (steps), wall-clock time (in minutes), average local batch sizes (bsz.), and top validation accuracy (acc.; in %)
-Schedule $H=16$ $H=4$ $H=1$
-steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
-Constant 2432 27.16 1024 92.31 2440 34.00 1024 92.65 2441 60.51 1024 91.11
-Constant 608 34.99 4096 63.50 608 37.10 4096 57.04 610 46.39 4096 40.40
-Constant 304 34.31 8192 12.51 304 35.34 8192 16.61 305 40.01 8192 14.78
-$\eta=0.8$ 576 90.73 4306 81.95 416 138.39 6001 75.09 241 50.67 10380 25.32
-$\eta=0.85$ 512 92.41 4900 81.95 912 141.51 2742 86.89 433 144.75 5767 72.66
-$\eta=0.9$ 2816 96.98 885 93.59 1982 126.38 1261 92.53 3940 111.92 634 93.75
+We examine the impact of local gradient steps $H$ (Table 1, Figure 1).
+According to Lin et al.
+(2020); Ortiz et al.
+(2021); Gu et al.
+(2023, 2024), local gradient methods may enhance generalization compared to synchronized ones $H=1$, although we use SHB instead of vanilla SGD in synchronized models.
+This is evident with a constant batch size of 8192 and 4 local gradient steps.
+For adaptive batch sizes, smaller $H$ values rapidly increase batch sizes (Figure 1), supporting our theory that a smaller $\eta$ with a larger $H$ ensures convergence rates of $\mathscr{O}(HM+\eta^{2})$.
+See Figure 1 and Section C.1 for more details.
+Table 1: Adaptive batch size strategies for ResNet-50 on CIFAR-10; no. of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best validation accuracy (acc.; in %)
+Schedule $H=32$ $H=16$ $H=4$ $H=1$
+steps time bsz. acc. steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
+Constant 1824 0.98 4096 67.02 1824 0.99 4096 75.32 1828 1.07 4096 88.12 1831 1.34 4096 89.40
+Constant 896 0.95 8192 44.27 912 0.98 8192 48.19 912 1.01 8192 78.81 915 1.15 8192 76.58
+Constant 576 1.07 12500 10.19 592 1.10 12500 20.89 596 1.13 12500 42.36 599 1.23 12500 53.80
+$\eta=0.8$ 928 1.13 7828 74.95 832 1.15 8906 76.50 744 1.16 10060 75.67 1241 1.41 6043 82.14
+$\eta=0.85$ 1088 1.18 7019 69.92 864 1.14 8607 75.32 756 1.16 9896 75.40 1270 1.43 5906 83.15
+$\eta=0.9$ 1216 1.15 6125 75.76 1088 1.16 6929 77.48 748 1.17 10022 74.35 1540 1.47 4868 84.61
+6.2 MicroLlama 300M on C4
+We pre-train MicroLlama 300M (Wang, 2024) using the C4 dataset (Raffel et al., 2020) processed by Allen AI [^1] See https://huggingface.co/datasets/allenai/c4, tokenized with Llama 2 7B’s tokenizer (Touvron et al., 2023) (vocabulary size: 32,000).
+Employing $M=4$ workers, the training budget is 4 billion tokens with a sequence length of 2048, and maximum batch sizes of 8192 (global) and 2048 (local) sequences, equating to $\approx$ 16M and 4M tokens.
+We use AdamW (Loshchilov and Hutter, 2019) as the inner optimizer, with a learning rate schedule featuring a 1% linear warmup and cosine decay, peaking at 0.001 and dropping to 0.0001.
+Table 2: Adaptive batch size strategies for MicroLlama 300M on C4; no.
+of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best validation loss (cross-entropy loss; estimated by 100 iterations)
+Schedule $H=32$ $H=16$ $H=4$
+steps time bsz. loss steps time bsz. loss steps time bsz. loss
+Constant 31744 10.59 512 4.10 15616 6.86 512 4.20 3888 11.91 512 3.93
+Constant 16384 10.53 1024 4.82 7936 10.64 1024 4.84 1968 11.31 1024 5.02
+Constant 8192 9.77 2048 5.72 4096 10.50 2048 5.73 992 10.96 2048 6.00
+$\eta=0.8$ 15360 11.13 1088 4.55 5632 10.96 1453 4.98 1216 11.13 1658 5.05
+$\eta=0.9$ 16384 11.54 1054 4.66 6400 11.22 1299 4.80 1360 11.18 1484 4.68
 Values of $\eta$.
-Table 3 shows that adaptive local batch size strategies can achieve comparable or superior validation accuracy to constant batch sizes with similar averages.
-For instance, with $H=16$ and $\eta=0.80$, the adaptive strategy reaches an average batch size akin to a constant batch size of 4096 but with nearly 50% higher accuracy and 32 fewer steps, illustrating the effectiveness of adaptive strategies in bridging the generalization gap in large-batch training.
-Figure 3: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$).
+Similar to the observation in Section 6.1, the adaptive batch size strategy generally achieves lower losses compared to their constant batch size counterparts, with a significant reduction in the number of training steps.
+It illustrates the effectiveness of our adaptive batch size strategy in bridging the generalization gap in large-batch training.
 Local gradient steps $H$.
-We first examine constant batch sizes (Table 3 and Figure 4).
-Local gradient methods, as noted in [50, 58, 24, 25], may improve generalization compared to synchronized methods ($H=1$), although we employ SHB in synchronized models instead of standard SGD.
-For adaptive local batch sizes, smaller $H$ values increase batch sizes more rapidly (Table 3), supporting our theory that a smaller $\eta$ should be used with a larger $H$ to maintain convergence rates of $\mathscr{O}(HM+\eta^{2})$.
-Refer to Figures 3 and B.2 for further illustrations.
-Figure 4: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies with $\eta\in\{0.25,0.375,0.5\}$ and $H\in\{16,4,1\}$.
-Number of workers $M$.
-We conducted additional experiments with $M=8$.
-Compared to the 4-worker scenario, increasing the number of workers reduces training time but lowers average accuracy, indicating poorer generalization.
-Nonetheless, the observations regarding local gradient steps $H$ and values of $\eta$ remain consistent with the 4-worker case.
-Figure 5: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies ($H=16$) with 8 workers.
-Table 3: Adaptive local batch size strategies for ResNet-50 on CIFAR-10 with 8 workers
-Schedule $H=16$ $H=4$ $H=1$
-steps time bsz. acc. steps time bsz. acc. steps time bsz. acc.
-Constant 2432 12.06 512 92.06 2440 15.45 512 92.87 2441 29.09 512 92.41
-Constant 1216 11.19 1024 88.88 1220 12.93 1024 87.87 1220 19.72 1024 86.80
-Constant 608 10.72 2048 66.94 608 11.59 2048 71.32 610 15.03 2048 59.57
-$\eta=0.8$ 213 11.55 5666 17.88 216 11.82 5655 15.98 218 12.29 5650 16.84
-$\eta=0.85$ 220 12.22 5610 19.47 222 12.30 5570 22.06 222 12.66 5569 21.99
-$\eta=0.9$ 261 15.25 5236 28.78 277 16.92 5052 33.07 278 18.26 4959 36.83
+Again, batch sizes grows more rapidly as the number of local gradient steps $H$ decreases, indicating an inverse scaling relationship between $\eta$ and $H$.
+This in turn leads to smaller batch sizes at the earlier stages of training and hence better generalization.
+We also observe from Figure 2(d) that, for $\eta=0.9$, the generalization gaps between different values of $H$ is narrower than that for $\eta=0.8$.
+(a) $H=32$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(b) $H=16$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(c) $H=4$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(d) Adaptive batch size strategies for $H\in\{32,16,4\}$
+Figure 2: Validation loss and local batch sizes of Local AdamW with adaptive batch size strategies for MicroLlama 300M on C4.
 7 Concluding Remarks
-This work introduces adaptive local batch size strategies for local gradient methods in data parallelism, focusing on reducing communication overheads by increasing local batch sizes independently.
-Although our analysis and experiments are limited to identical data settings, the algorithm is applicable to heterogeneous scenarios.
-Future work will explore convergence in diverse settings and federated learning [54, 33], including personalized approaches [27].
-Extensions to asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75] are also promising for improving communication efficiency and addressing worker heterogeneity.
+This work presents adaptive batch size strategies for local gradient methods under data parallelism to reduce communication overhead by increasing local batch sizes.
+To address possible efficiency issues in practical implementation, we propose an approximation of the norm test under the context of local gradient methods, by taking advantage of the availability of local batch gradients.
+Illustrated with experiments in both image classification and language modeling task, our proposed approach is able to reduce the generalization gap of large-batch training while only adding a small amount of additional training time.
+Future research will examine convergence in varied settings and federated learning (McMahan et al., 2017; Kairouz et al., 2021), including personalized approaches (Hanzely et al., 2023).
+Asynchronous SGD (Mishchenko et al., 2022; Koloskova et al., 2022; Nguyen et al., 2022; Leconte et al., 2024; Islamov et al., 2024) and hierarchical SGD (Lin et al., 2020; Castiglia et al., 2021; Wang et al., 2022) extensions are promising for enhancing communication efficiency and tackling worker heterogeneity.
 Acknowledgments
+The research of Tim Tsz-Kit Lau is supported in part through the computational resources and staff contributions from the Data Science Institute at the University of Chicago, through its AI + Science Research Funding Initiatives.
 The research of Han Liu is supported by NIH R01LM01372201, NSF RI 1840857, NSF TRIPOD 1740735, NSF DMS1454377-CAREER, NSF IIS 1546482, along with an Alfred P.
@@
 The research of Mladen Kolar is supported in part by NSF ECCS-2216912.
-This research is supported in part through the computational resources and staff contributions from the Data Science Institute at the University of Chicago, through its AI + Science Research Funding Initiatives.
 Appendix
-Appendix A Proofs of Main Text
-We present the omitted proofs of the main text, mainly by simplifying the technique in [69].
+Appendix A Main Algorithm
+Algorithm A.1 Adaptive Local Batch Size Schedules for Local Gradient Methods
+no.
+of workers $M\in\mathbb{N}^{*}$, no.
+of communication steps $K\in\mathbb{N}^{*}$, no.
+of local steps $H\in\mathbb{N}^{*}$, no.
+of training samples $N\in\mathbb{N}^{*}$, global step counter $k=0$, processed samples counter $B=0$, training sets $(\mathcal{D}_{m})_{m\in\llbracket M\rrbracket}=(\{\xi_{i}^{m}\}_{i=1}^{n_{m}})_{m\in\llbracket M\rrbracket}$, initial model parameters $x_{0,0}^{m}=x_{0}\in\mathbb{R}^{d}$for all $m\in\llbracket M\rrbracket$, initial local batch sizes $(b_{0,0}^{m})_{m\in\llbracket M\rrbracket}$, learning rate schedules $(\alpha_{k,h})_{k\in\llbracket 0,K-1\rrbracket,h\in\llbracket 0,H-1\rrbracket}$, $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$$B<N$worker $m=1,\ldots,M$in parallel local step $h=0,\ldots,H-1$
+Sample the local batch $\mathcal{B}_{k,h}^{m}$ of size $b_{k,h}^{m}\coloneqq|\mathcal{B}_{k,h}^{m}|$ uniformly from $\mathcal{D}_{m}$
+Compute per-sample gradients $\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$ for each $i\in\mathcal{B}_{k,h}^{m}$
+Compute the local batch gradient $\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})=\frac{1}{b_{k,h}^{m}}\sum_{i\in\mathcal{B}_{k,h}^{m}}\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$
+Compute $\mathrm{Var}_{i\in\mathcal{B}_{k,h}^{m}}(\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m}))$ by applying (5)
+Compute the local norm test statistic $\mathsf{T}_{k,h}^{m}\coloneqq\left\lceil\left(\mathrm{Var}_{i\in\mathcal{B}_{k,h}^{m}}(\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m}))\right)\Big{/}\left(\eta^{2}\|\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})\|^{2}\right)\right\rceil$
+Determine the next local batch sizes: $b_{k,h+1}^{m}=\max\{\mathsf{T}_{k,h}^{m},b_{k,h}^{m}\}$
+[…]
+Update the global model: $x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}$ for each $m\in\llbracket M\rrbracket$triangleright all-reduce
+$B\leftarrow B+\sum_{m=1}^{M}\sum_{h=0}^{H-1}b_{k,h}^{m}$
+[…]
+[…]
+[…]
+[…]
+[…]
+Algorithm A.2 Adaptive Batch Size Schedules for Local Gradient Methods (Actual Implementation)
+no.
+of workers $M\in\mathbb{N}^{*}$, no.
+of communication steps $K\in\mathbb{N}^{*}$, no.
+of local steps $H\in\mathbb{N}^{*}$, no.
+of training samples $N\in\mathbb{N}^{*}$, global step counter $k=0$, processed samples counter $B=0$, training sets $(\mathcal{D}_{m})_{m\in\llbracket M\rrbracket}=(\{\xi_{i}^{m}\}_{i=1}^{n_{m}})_{m\in\llbracket M\rrbracket}$, initial model parameters $x_{0,0}^{m}=x_{0}\in\mathbb{R}^{d}$for all $m\in\llbracket M\rrbracket$, initial local batch sizes $(b_{0}^{m})_{m\in\llbracket M\rrbracket}$, learning rate schedules $(\alpha_{k,h})_{k\in\llbracket 0,K-1\rrbracket,h\in\llbracket 0,H-1\rrbracket}$, $(\eta_{m})_{m\in\llbracket M\rrbracket}\subset(0,1)$$B<N$worker $m=1,\ldots,M$in parallel local step $h=0,\ldots,H-1$
+Sample the local batch $\mathcal{B}_{k,h}^{m}$ of size $b_{k}^{m}\coloneqq|\mathcal{B}_{k,h}^{m}|$ uniformly from $\mathcal{D}_{m}$
+Compute the local batch gradient $\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})=\frac{1}{b_{k,h}^{m}}\sum_{i\in\mathcal{B}_{k,h}^{m}}\nabla f_{m}(x_{k,h}^{m};\xi_{i}^{m})$
+[…]
+Update the global model: $x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}$ for each $m\in\llbracket M\rrbracket$triangleright all-reduce
+Compute the averaged batch gradient of all workers: $\overline{g}_{k,H}\coloneqq\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k,H}^{m}}(x_{k,H}^{m})$triangleright all-reduce
+Compute $\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k}^{m};\xi_{i}^{m})\right)=b_{k}^{m}\cdot\frac{1}{M-1}\sum_{m=1}^{M}\left\lVert\nabla F_{\mathcal{B}_{k,H}^{m}}(x_{k,H}^{m})-\overline{g}_{k,H}\right\rVert^{2}$
+Compute the norm test statistic $\mathsf{T}_{k}^{m}\coloneqq\left\lceil\left(\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k,H}^{m};\xi_{i}^{m})\right)\right)\Big{/}\left(M\eta^{2}\left\lVert\overline{g}_{k,H}\right\rVert^{2}\right)\right\rceil$
+Determine the next local batch sizes: $b_{k+1}^{m}=\max\{\mathsf{T}_{k}^{m},b_{k}^{m}\}$
+$B\leftarrow B+HMb_{k}^{m}$
+[…]
+[…]
+[…]
+[…]
+[…]
+Appendix B Proofs of Main Text
+We present the omitted proofs of the main text, mainly by simplifying the technique in Stich and Karimireddy (2020).
 Notation.
-Let us define
-$(\forall k\in\mathbb{N})\quad\overline{x}_{k+1}\coloneqq\overline{x}_{k}-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}g_{k}^{m},$ (A.1)
-where $g_{k}^{m}\coloneqq\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$ and $\overline{x}_{0}\coloneqq x_{0}$.
+Let us define a sequence of (virtual) averaged iterates $(\overline{x}_{k})_{k\in\mathbb{N}}$, given by
+$(\forall k\in\mathbb{N})\quad\overline{x}_{k+1}\coloneqq\overline{x}_{k}-\alpha_{k}\overline{g}_{k},$ (B.1)
+where $\overline{g}_{k}\coloneqq\frac{1}{M}\sum_{m=1}^{M}g_{k}^{m}$ with $g_{k}^{m}\coloneqq\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})$ and $\overline{x}_{0}\coloneqq x_{0}$.
 Let use recall that $\mathbb{E}_{k}\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$, $\mathbb{E}_{\mathcal{F}_{k}^{m}}\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}^{m}]$, and $\mathbb{E}$ denotes the unconditional expectation.
 We also define the notion of slowly decreasing learning rates, called $\tau$-slow sequences.
-Definition A.1 ($\tau$-slow sequence; Definition 10 in in [69]).
+Definition B.1 ($\tau$-slow sequence; Definition 10 in Stich and Karimireddy, 2020).
 A sequence $(a_{k})_{k\in\mathbb{N}}\subset(0,\infty)$ with positive values is $\tau$-slow decreasing for $\tau\geqslant 1$ if
@@
 The sequence $(a_{k})_{k\in\mathbb{N}}$ is $\tau$-slow increasing if $(1/a_{k})_{k\in\mathbb{N}}$ is $\tau$-slow decreasing.
-A.1 Technical Lemmas
+B.1 Technical Lemmas
 We first state without proofs some preparatory lemmas, which are either standard results or whose proofs are be found in the cited work.
-Lemma A.1.
+Lemma B.1.
 For any $M$ vectors $v_{1},\ldots,v_{M}\in\mathbb{R}^{d}$, we have
-$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert\leqslant\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert.$ (A.2)
-Lemma A.2.
+$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert\leqslant\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert.$ (B.2)
+Lemma B.2.
 For any $M$ vectors $v_{1},\ldots,v_{M}\in\mathbb{R}^{d}$, we have
-$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert^{2}\leqslant M\sum_{m=1}^{M}\left\lVert w_{m}\right\rVert^{2}.$ (A.3)
-Lemma A.3.
+$\left\lVert\sum_{m=1}^{M}v_{m}\right\rVert^{2}\leqslant M\sum_{m=1}^{M}\left\lVert v_{m}\right\rVert^{2}.$ (B.3)
+Lemma B.3.
 For any two vectors $x,y\in\mathbb{R}^{d}$ and any positive number $\rho>0$, we have
-$\left\lVert x+y\right\rVert^{2}\leqslant(1+\rho)\left\lVert a\right\rVert^{2}+(1+1/\rho)\left\lVert b\right\rVert^{2}.$ (A.4)
-Lemma A.4 (Cauchy–Schwarz inequality).
+$\left\lVert x+y\right\rVert^{2}\leqslant(1+\rho)\left\lVert a\right\rVert^{2}+(1+1/\rho)\left\lVert b\right\rVert^{2}.$ (B.4)
+Lemma B.4 (Cauchy–Schwarz inequality).
 For any two vectors $x,y\in\mathbb{R}^{d}$ and any positive number $\rho>0$, we have
-$2\langle x,y\rangle\leqslant\rho\left\lVert a\right\rVert^{2}+\rho^{-1}\left\lVert b\right\rVert^{2}.$ (A.5)
-Lemma A.5.
+$2\langle x,y\rangle\leqslant\rho\left\lVert a\right\rVert^{2}+\rho^{-1}\left\lVert b\right\rVert^{2}.$ (B.5)
+Lemma B.5.
 If $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is $L$-Lipschitz smooth for some $L>0$ (Assumption 1), then we have
-$(\forall x\in\mathbb{R}^{d})\quad\left\lVert\nabla F(x)\right\rVert^{2}\leqslant 2L(F(x)-F^{\star}).$ (A.6)
-Lemma A.6 (Lemma 2 in [68]; Lemma 13 in [69]).
+$(\forall x\in\mathbb{R}^{d})\quad\left\lVert\nabla F(x)\right\rVert^{2}\leqslant 2L(F(x)-F^{\star}).$ (B.6)
+Lemma B.6 (Lemma 2 in Stich, 2019b; Lemma 13 in Stich and Karimireddy, 2020).
 For any nonnegative sequence $(u_{k})_{k\in\mathbb{N}}$ and parameters $\kappa\geqslant\chi>0$, $\sigma\geqslant 0$ and $K\in\mathbb{N}$, there exists a constant $\alpha\leqslant 1/\kappa$ such that for constant learning rates $(\alpha_{k}\equiv\alpha)_{k\in\mathbb{N}}$ and weights $w_{k}\coloneqq(1-\chi\alpha)^{-k-1}$, we have
@@
 where $W_{K}\coloneqq\sum_{k=1}^{K}w_{k}$.
-Lemma A.7 (Lemma 14 in [69]).
+Lemma B.7 (Lemma 14 in Stich and Karimireddy, 2020).
 For any nonnegative sequence $(u_{k})_{k\in\mathbb{N}}$ and parameters $\kappa\geqslant 0$, $\sigma\geqslant 0$ and $K\in\mathbb{N}$, there exists a constant $\alpha\leqslant 1/\kappa$ such that for constant learning rates $(\alpha_{k}\equiv\alpha)_{k\in\mathbb{N}}$, we have
 $\Psi_{K}\coloneqq\frac{1}{K+1}\sum_{k=0}^{K}\left(\frac{u_{k}}{\alpha_{k}}-\frac{u_{k+1}}{\alpha_{k}}+\sigma\alpha_{k}\right)\leqslant\frac{\kappa u_{0}}{K+1}+2\sqrt{\frac{\sigma u_{0}}{K+1}}.$
-A.2 Proof of Theorems 1, 2 and 3
+B.2 Proofs of Theorems 1, 2 and 3
 We first prove a descent lemma for ($\mu$-strongly) convex objectives.
-Lemma A.8 (Descent lemma for ($\mu$-strongly) convex objectives).
+Lemma B.8 (Descent lemma for ($\mu$-strongly) convex objectives).
 Suppose that Assumptions 1 and 2 hold with $\mu>0$.
 Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (9) with $\eta_{m}\equiv\eta\in(0,1)$.
-If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
+If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(4L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence $(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (B.1) satisfies
 $\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\leqslant\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{\alpha_{k}}{2M}\sum_{m=1}^{M}\mathbb{E}[F(x_{k}^{m})-F^{\star}]+\frac{3L\alpha_{k}}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
 Proof.
-The proof is similar to that of Lemma 25 in [69]. We state it here for convenience and completeness. Expanding we have
+The proof is similar to that of Lemma 25 in Stich and Karimireddy (2020). We state it here for convenience and completeness. Expanding we have
 $\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}=\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle g_{k}^{m},x_{k}^{m}-x^{\star}\rangle+\frac{\alpha_{k}^{2}}{M^{2}}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}+\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle g_{k}^{m},x_{k}^{m}-\overline{x}_{k}\rangle.$
@@
 $\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$ $=\left\lVert\sum_{m=1}^{M}\nabla F(x_{k}^{m})\right\rVert^{2}+\sum_{m=1}^{M}\mathbb{E}\left\lVert g_{k}^{m}-\nabla F(x_{k}^{m})\right\rVert^{2}$
-$\leqslant M\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\eta^{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.3) and (8) (A.7)
-[…]
+$\leqslant M\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\eta^{2}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (B.3) and (12) (B.7)
+$\leqslant 2L(M+\eta^{2})\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)$ by (B.6) (B.8)
 On the other hand, by Assumption 2, we also have
-$-2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle\leqslant-\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}-2\left(F(x_{k}^{m})-F^{\star}\right).$ (A.8)
-By Cauchy–Schwarz inequality (A.5) with $\rho=2L$, we have
+$-2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle\leqslant-\mu\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}-2\left(F(x_{k}^{m})-F^{\star}\right).$ (B.9)
+By Cauchy–Schwarz inequality (B.5) with $\rho=2L$, we have
 $2\langle\nabla F(x_{k}^{m}),x_{k}^{m}-\overline{x}_{k}\rangle$ $\leqslant 2L\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+\frac{1}{2L}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$
-[…]
-By (A.4), we also have
-$-\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}\leqslant-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$ (A.10)
-Thus, plugging in (A.6), (A.9) and (A.10), we obtain
+$\leqslant 2L\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}+F(x_{k}^{m})-F^{\star}$ by (B.6) (B.10)
+By (B.4), we also have
+$-\left\lVert x_{k}^{m}-x^{\star}\right\rVert^{2}\leqslant-\frac{1}{2}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}+\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$ (B.11)
+Thus, plugging in (B.8), (B.10) and (B.11), we obtain
 $\mathbb{E}_{k}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}$ $\leqslant\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{2\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m}),x_{k}^{m}-x^{\star}\rangle+\frac{2L\alpha_{k}^{2}(M+\eta^{2})}{M^{2}}\sum_{m=1}^{M}\left(F(x_{k}^{m})-F^{\star}\right)$
@@
 We also prove a descent lemma for general nonconvex objectives.
-Lemma A.9 (Descent lemma for nonconvex objectives).
+Lemma B.9 (Descent lemma for nonconvex objectives).
 Suppose that Assumption 1 holds.
 Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (9) with $\eta_{m}\equiv\eta\in(0,1)$.
-If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence$(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (A.1) satisfies
+If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant M/(2L(M+\eta^{2}))$ for all $k\in\mathbb{N}$, then the sequence $(\overline{x}_{k})_{k\in\mathbb{N}}$ defined in (B.1) satisfies
 $\mathbb{E}[F(\overline{x}_{k+1})]\leqslant\mathbb{E}[F(\overline{x}_{k})]-\frac{\alpha_{k}}{4M}\sum_{m=1}^{M}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{L^{2}\alpha_{k}}{2M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
 Proof.
-The proof is similar to that of Lemma 26 in [69]. We state it here for convenience and completeness.
+The proof is similar to that of Lemma 26 in Stich and Karimireddy (2020). We state it here for convenience and completeness.
 By Assumption 1, we have
@@
 $\mathbb{E}_{k}[F(\overline{x}_{k+1})]$ $\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}}{2M^{2}}\mathbb{E}_{k}\left\lVert\sum_{m=1}^{M}g_{k}^{m}\right\rVert^{2}$
-$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (A.7)
-$=F(\overline{x}_{k})-\left(\frac{\alpha_{k}}{M}-\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$
-$\qquad+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle.$
-By Cauchy–Schwarz inequality (A.5) with $\rho=1$ and Assumption 1, we have
+$\leqslant F(\overline{x}_{k})-\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle+\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ by (B.7)
+$=F(\overline{x}_{k})-\left(\frac{\alpha_{k}}{M}-\frac{L\alpha_{k}^{2}(M+\eta^{2})}{2M^{2}}\right)\sum_{m=1}^{M}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}+\frac{\alpha_{k}}{M}\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle.$
+By Cauchy–Schwarz inequality (B.5) with $\rho=1$ and Assumption 1, we have
 $\sum_{m=1}^{M}\langle\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k}),\nabla F(x_{k}^{m})\rangle$ $\leqslant\frac{1}{2}\sum_{m=1}^{M}\left[\left\lVert\nabla F(x_{k}^{m})-\nabla F(\overline{x}_{k})\right\rVert{\nabla F(x_{k}^{m})}^{2}+\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}\right]$
@@
 We then derive an upper bound for $\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$.
-Lemma A.10.
-Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (8) with $\eta_{m}\equiv\eta\in(0,1)$, and $(\overline{x}_{k})_{k\in\mathbb{N}}$ be the sequence defined in (A.1).
+Lemma B.10.
+Let $(x_{k}^{m})_{k\in\mathbb{N},m\in\llbracket M\rrbracket}$ be the sequence of the iterates of Local SGD (2) with the (exact variance) local norm test (9) with $\eta_{m}\equiv\eta\in(0,1)$, and $(\overline{x}_{k})_{k\in\mathbb{N}}$ be the sequence defined in (B.1).
 If the sequence of learning rates $(\alpha_{k})_{k\in\mathbb{N}}$ satisfies $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$ and $(\alpha_{k}^{2})_{k\in\mathbb{N}}$ is $H$-slow decreasing, then we have
-$\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{10LHM}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.11)
+$\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{1}{10LHM}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (B.12)
 Moreover, for any $H$-slow increasing nonnegative sequence $(w_{k})_{k\in\mathbb{N}}$ we have
@@
 $\leqslant\frac{3H}{2M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
-$\qquad\qquad+\frac{3}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right\rVert^{2}$ by (A.4) and (A.3)
+$\qquad\qquad+\frac{3}{M}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert g_{k-h}^{m}-\nabla F(x_{k-h}^{m})\right\rVert^{2}$ by (B.4) and (B.3)
 $\leqslant\frac{3(H+\eta^{2})}{2K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\alpha_{k-h}^{2}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$
@@
 Then, we can simply the inequality:
-$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{3(H+\eta^{2})\alpha_{k}^{2}}{K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.12)
+$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{3(H+\eta^{2})\alpha_{k}^{2}}{K}\sum_{m=1}^{M}\sum_{h=0}^{H-1}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (B.13)
 Since $\alpha_{k}\leqslant 1/(10L(HM+\eta^{2}))$ for all $k\in\mathbb{N}$, we have
@@
 $\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}\frac{1}{H}\sum_{h=0}^{H-1}w_{k-h}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}$
-$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (A.13)
+$\leqslant\frac{1}{5LM}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k-h}^{m})\right\rVert^{2}.$ (B.14)
 Proof of Theorem 1.
-Using Lemma A.10 with $\mu>0$ and (A.6), we have
-$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}].$ (A.14)
-Since $\alpha_{k}\equiv\alpha\leqslant 1/(10L(HM+\eta^{2}))\leqslant M/(4L(M+\eta^{2}))$, by Lemma A.8, we have
+Using Lemma B.10 with $\mu>0$ and (B.6), we have
+$\frac{3L}{M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}\leqslant\frac{2}{5M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k-h}^{m})-F^{\star}].$ (B.15)
+Since $\alpha_{k}\equiv\alpha\leqslant 1/(10L(HM+\eta^{2}))\leqslant M/(4L(M+\eta^{2}))$, by Lemma B.8, we have
 $\frac{1}{2M}\sum_{m=1}^{M}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{1}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{1}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}+\frac{3L}{M}\sum_{m=1}^{M}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}.$
-Multiplying the weight $w_{k}$ and summing over $k\in\llbracket 0,K\rrbracket$, and by (A.14), we have
+Multiplying the weight $w_{k}$ and summing over $k\in\llbracket 0,K\rrbracket$, and by (B.15), we have
 $\quad\,\frac{1}{2M}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]$
@@
 which is equivalent to
-$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{10}{W_{K}}\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right).$ (A.15)
+$\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}[F(x_{k}^{m})-F^{\star}]\leqslant\frac{10}{W_{K}}\sum_{k=0}^{K}\left(\frac{w_{k}}{\alpha_{k}}\left(1-\frac{\mu\alpha_{k}}{2}\right)\mathbb{E}\left\lVert\overline{x}_{k}-x^{\star}\right\rVert^{2}-\frac{w_{k}}{\alpha_{k}}\mathbb{E}\left\lVert\overline{x}_{k+1}-x^{\star}\right\rVert^{2}\right).$ (B.16)
 We observe that $1-\mu\alpha/2\geqslant 1-\mu/(20L(HM+\eta^{2}))\geqslant 1-1/(8HM)$, we choose $w_{k}=(1-\mu\alpha/2)^{-k-1}$ which are $2H$-slow increasing.
-Then by Lemma A.6, we obtain the desired result.
+Then by Lemma B.6, we obtain the desired result.
 ∎
 Proof of Theorem 2.
-For the convex case with $\mu=0$, we simply use Lemma A.6 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$. ∎
+For the convex case with $\mu=0$, we simply use Lemma B.6 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$. ∎
 Proof of Theorem 3.
-By Lemma A.9 and (A.6), and define $u_{k}\coloneqq 4\cdot\mathbb{E}[F(\overline{x}_{k})-F^{\star}]$, we have
+By Lemma B.9 and (B.6), and define $u_{k}\coloneqq 4\cdot\mathbb{E}[F(\overline{x}_{k})-F^{\star}]$, we have
 $\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}$ $\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right)+\frac{2L\alpha_{k}}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert x_{k}^{m}-\overline{x}_{k}\right\rVert^{2}$
 $\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right)+\frac{2\alpha_{k}}{15LMW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2},$
-by (A.13). We simplify this to
+by (B.14). We simplify this to
 $\left(1-\frac{2\alpha_{k}}{15L}\right)\frac{1}{MW_{K}}\sum_{m=1}^{M}\sum_{k=0}^{K}w_{k}\mathbb{E}\left\lVert\nabla F(x_{k}^{m})\right\rVert^{2}\leqslant\frac{1}{W_{K}}\sum_{k=0}^{K}w_{k}\left(\frac{u_{k}}{4\alpha_{k}}-\frac{u_{k+1}}{4\alpha_{k}}\right).$
-Applying Lemma A.7 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$ gives the desired result.
-Appendix B Details of Numerical Experiments
-All experiments are conducted on a Slurm cluster.
-For experiments with 4 GPUs, each node has either 4 NVIDIA GEFORCE RTX 2080 Ti (11GB) GPUs or 4 NVIDIA A40 (48GB) GPUs or 4 NVIDIA L40s (48GB) GPUs.
-For experiments with 8 GPUs, each node has 8 NVIDIA A100 (80GB) GPUs.
-For CPUs, each node has either 64 cores or 96 cores.
-The total RAM is 512GB or 1TB.
-B.1 Three-layer Convolutional Neural Network on MNIST
-Table 4: Training hyperparameters for three-layer CNN on MNIST
-Model Three-layer CNN on MNIST
-Training samples 6M (100 epochs)
-Weight initialization Default
-Optimizer Adam
-Learning rate schedule None
-$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
-$\varepsilon$ $10^{-8}$
-Base learning rate 0.0008
-Base batch size 8
-Maximum batch size 50000
-Weight decay 0
-Base global batch size 4096
-Base local batch size 1024
-Maximum global batch size 60000
-Maximum local batch size 15000
-Weight decay 0
-Figure 6: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies with three different seeds.
-Figure 7: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
-Figure 8: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
-Figure 9: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
-B.2 ResNet-50 on CIFAR-10
+Applying Lemma B.7 with $w_{k}\equiv 1$ for all $k\in\mathbb{N}$ gives the desired result.
+Appendix C Details of Numerical Experiments
+For experiments with 4 GPUs, each node has either NVIDIA L40S (48GB; for ResNet-50 on CIFAR-10) or A40 (48GB; for MicroLlama 300M on C4) or A100 (SXM4 80GB; for ResNet-101 on ImageNet) GPUs.
+C.1 ResNet-50 on CIFAR-10
 Table 3: Training hyperparameters for ResNet-50 on CIFAR-10
 Model ResNet-50 on CIFAR-10
-Training samples 10M (200 epochs)
+Training samples 30,000,000 (600 epochs)
 Weight initialization Default
@@
 Learning rate schedule Linear warmup + cosine decay
-Learning rate warmup (samples) 1M (10%)
+Learning rate warmup (samples) 1,000,000 (10%)
 Peak learning rate 0.05
-Minimum learning rate 0.005
-Base learning rate 0.05
+Base learning rate 0.005
+Learning rate scaling rule linear (for constant batch sizes)
 Base global batch size 256
 Base local batch size 64
-Maximum global batch size 50000
-Maximum local batch size 12500
-Weight decay 0.0005
+Maximum global batch size 50,000
+Maximum local batch size 12,500
+Weight decay 0.0001
 Momentum 0.9
-Figure 10: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with three different seeds.
-Figure 11: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with two other different seeds.
-Figure 12: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with three different seeds.
-Figure 13: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with three different seeds.
-Figure 14: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies with 8 workers with three different seeds.
-Figure 15: Validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies (local gradient steps $H=16$) with 8 workers with two other different seeds.
-Figure 16: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=4$) with 8 workers with three different seeds.
-Figure 17: Validation accuracy and local batch sizes curves of Local Adam with adaptive batch size strategies (local gradient steps $H=1$) with 8 workers with three different seeds.
+Precision bfloat16
+Data-parallel size 4
+(a) $H=1$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+Figure 3: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies for ResNet-50 on CIFAR-10.
+Table 4: Adaptive batch size strategies for ResNet-50 on CIFAR-10; mean (standard deviation) of no.
+of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best validation accuracy (acc.; in %) of all three seeds
+Schedule $H=32$ $H=16$
+steps time bsz. acc. steps time bsz. acc.
+Constant 1824 0.98 (0.00) 4096 75.23 (7.19) 1824 0.99 (0.00) 4096 79.62 (4.39)
+Constant 896 0.95 (0.00) 8192 33.92 (10.78) 912 0.98 (0.00) 8192 31.23 (14.77)
+Constant 576 1.07 (0.00) 12500 10.70 (0.79) 592 1.10 (0.00) 12500 14.92 (5.49)
+$\eta=0.8$ 950 (98) 1.14 (0.02) 7787 (607) 71.49 (3.09) 827 (25) 1.14 (0.01) 9030 (209) 72.15 (3.82)
+$\eta=0.85$ 1088 (0) 1.16 (0.02) 6943 (129) 70.54 (1.03) 880 (43) 1.14 (0.01) 8457 (374) 72.41 (2.52)
+$\eta=0.9$ 1163 (67) 1.14 (0.01) 6346 (341) 73.02 (2.82) 1072 (43) 1.15 (0.00) 7027 (238) 73.22 (3.72)
+Schedule $H=4$ $H=1$
+steps time bsz. acc. steps time bsz. acc.
+Constant 1828 1.06 (0.00) 4096 88.72 (0.52) 1831 1.34 (0.00) 4096 89.31 (0.34)
+Constant 912 1.01 (0.00) 8192 76.56 (2.29) 915 1.15 (0.00) 8192 78.19 (2.63)
+Constant 596 1.14 (0.00) 12500 35.70 (11.59) 599 1.23 (0.00) 12500 40.60 (15.22)
+$\eta=0.8$ 750 (25) 1.16 (0.00) 10001 (346) 74.66 (0.88) 1023 (211) 1.36 (0.05) 7548 (1554) 80.56 (2.08)
+$\eta=0.85$ 771 (41) 1.16 (0.01) 9736 (494) 75.57 (0.21) 935 (297) 1.35 (0.08) 8525 (2399) 78.89 (4.03)
+$\eta=0.9$ 756 (14) 1.16 (0.01) 9916 (166) 74.77 (0.42) 1246 (354) 1.42 (0.07) 6415 (2094) 81.73 (3.65)
+Observations.
+Note that due to the stochastic nature of drawing data samples when computing gradients, the adaptive batch size strategies might give rise to different adaptive batch size schedules when different random seeds.
+We hence perform the training runs for three different random seeds to evaluate the robustness of the proposed strategies (see results of the other two seeds in Figures 4 and 5).
+We observe from Table 4 that, as the number of local steps $H$ grows, constant batch sizes might have more varied performance (in terms of standard deviation of best validation accuracy) than the proposed strategies.
+Also notice that, due to the use of relatively large batch sizes, all the experiments make use of the technique of gradient accumulation, which essentially consists of serial operations using for-loops.
+Therefore, the benefit of using large batch sizes is not reflected in the total training time.
+If a larger number of GPUs or stronger GPUs with more memory are used, training with large batch sizes will consume substantially shorter time.
+Consequently, it is better to measure the computational efficiency of different training runs using the number of total training steps (Shallue et al., 2019).
+(a) $H=32$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(b) $H=16$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(c) $H=4$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(d) $H=1$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(e) Adaptive batch size strategies for $H\in\{32,16,4\}$
+Figure 4: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies for ResNet-50 on CIFAR-10 of the second seed.
+(a) $H=32$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(b) $H=16$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(c) $H=4$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(d) $H=1$ with $b_{k}^{m}\in\{4096,8192,12500\}$ and $\eta\in\{0.8,0.85,0.9\}$
+(e) Adaptive batch size strategies for $H\in\{32,16,4\}$
+Figure 5: Validation accuracy and local batch sizes of Local SHB with adaptive batch size strategies for ResNet-50 on CIFAR-10 of the third seed.
+C.2 MicroLlama 300M on C4
+Table 5: Training hyperparameters for MicroLlama 300M on C4
+Model MicroLlama 300M on C4
+Training samples (sequences) 2M
+Sequence length 2048 tokens
+Weight initialization Default
+Normalization Layer Normalization
+Optimizer AdamW
+$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
+Learning rate schedule Linear warmup + cosine decay
+Learning rate warmup (samples) 20,000 (1%)
+Peak learning rate 0.001
+Base learning rate 0.0001
+Learning rate scaling rule linear (for constant batch sizes)
+Base global batch size 256
+Base local batch size 64
+Maximum global batch size 8,192
+Maximum local batch size 2,048
+Weight decay 0.1
+Weight decay skip bias Yes
+Precision bfloat16
+Dropout 00
+Gradient clipping $1.0$
+Data-parallel size 4
+Table 6: Adaptive batch size strategies for MicroLlama 300M on C4; mean (standard deviation) of no.
+of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best validation loss (cross-entropy loss; estimated by 100 iterations) of all three seeds
+Schedule $H=32$ $H=16$
+steps time bsz. loss steps time bsz. loss
+Constant 31744 10.55 (0.04) 512 4.20 (0.18) 16384 10.51 (0.03) 512 5.88 (0.21)
+Constant 16384 10.51 (0.03) 1024 4.96 (0.21) 7936 10.60 (0.60) 1024 5.08 (0.21)
+Constant 8192 9.77 (0.01) 2048 5.88 (0.21) 4096 10.47 (0.02) 2048 5.84 (0.10)
+$\eta=0.8$ 14336 (1774) 11.00 (0.12) 1162 (155) 4.81 (0.31) 6059 (391) 11.02 (0.05) 1356 (89) 4.96 (0.18)
+$\eta=0.9$ 13995 (2132) 11.15 (0.44) 1207 (134) 4.88 (0.19) 7168 (678) 11.11 (0.36) 1163 (130) 4.72 (0.19)
+Schedule $H=4$
+steps time bsz. loss
+Constant 14336 11.00 (0.12) 512 4.81 (0.31)
+Constant 1968 11.28 (0.04) 1024 5.14 (0.13)
+Constant 992 10.92 (0.03) 2048 5.99 (0.04)
+$\eta=0.8$ 1206 (177) 11.14 (0.02) 1701 (260) 5.15 (0.38)
+$\eta=0.9$ 1238 (199) 11.14 (0.06) 1668 (303) 4.94 (0.33)
+Observations.
+For the task of language modeling with MicroLlama 300M on the C4 dataset, we also perform the training runs with three different random seeds.
+Similar conclusions to those in the main text can also be made, even though the proposed adaptive batch size strategies lead to different batch size growth patterns when using different random seeds (see Figures 6 and 7).
+(a) $H=32$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(b) $H=16$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(c) $H=4$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(d) Adaptive batch size strategies for $H\in\{32,16,4\}$
+Figure 6: Validation loss and local batch sizes of Local AdamW with adaptive batch size strategies for MicroLlama 300M on C4 of the second seed.
+(a) $H=32$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(b) $H=16$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(c) $H=4$ with $b_{k}^{m}\in\{512,1024,2048\}$ and $\eta\in\{0.8,0.9\}$
+(d) Adaptive batch size strategies for $H\in\{32,16,4\}$
+Figure 7: Validation loss and local batch sizes of Local AdamW with adaptive batch size strategies for MicroLlama 300M on C4 of the third seed.
+C.3 ResNet-101 on ImageNet
+We also train a ResNet-101 (He et al., 2016) on the ImageNet dataset (Russakovsky et al., 2015) to evaluate our method on models and datasets of larger scales.
+Table 7: Training hyperparameters for ResNet-101 on ImageNet
+Model ResNet-101 on ImageNet
+Training samples 256,233,400 (200 epochs)
+Weight initialization Default
+Optimizer Momentum SGD (SHB)
+Learning rate schedule Linear warmup + cosine decay
+Learning rate warmup (samples) 6,405,835 (2.5%)
+Peak learning rate 2.5
+Base learning rate 0.25
+Base global batch size 512
+Base local batch size 128
+Maximum global batch size 52,000
+Maximum local batch size 13,000
+Weight decay 0.0001
+Momentum 0.9
+Precision bfloat16
+Data-parallel size 4
+Table 8: Adaptive batch size strategies for ResNet-101 on ImageNet; no.
+of total steps (steps), wall-clock time (in hours), average local batch sizes (bsz.), and best top-1 validation accuracy (acc.; in %), and best top-5 validation accuracy (acc.@5; in %)
+Schedule $H=32$ $H=16$
+steps time bsz. acc. acc.@5 steps time bsz. acc. acc.@5
+Constant 10656 14.56 6000 59.20 81.84 10672 14.78 6000 63.76 85.18
+Constant 4896 14.35 13000 38.77 63.30 4912 14.34 13000 50.87 74.89
+$\eta=0.9$ 5216 14.53 12284 50.61 74.59 5072 14.64 12603 55.63 78.86
+$\eta=0.95$ 5280 14.31 12124 49.13 73.23 5088 15.09 12573 58.41 81.17
+Schedule $H=4$
+steps time bsz. acc. acc.@5
+Constant 10676 17.20 6000 71.28 89.97
+Constant 4924 15.41 13000 62.66 84.33
+$\eta=0.9$ 4952 15.62 12931 65.90 86.47
+$\eta=0.95$ 4976 16.75 12873 67.05 87.24
+Observations.
+We observe from Table 8 that the proposed adaptive batch size strategy with $\eta=0.9$ achieves better generalization performance (in terms of both accuracy and top-5 accuracy) than that of a constant large local batch size of 13,000 but worse than that of a constant small local batch size of 6,000 (i.e., generalization gap), while requiring also an intermediate amount of training time.
+The performance gap grows with the number of local steps $H$ (cf.
+Figures 8, 9 and 10).
+Figure 8: Validation accuracy, top-5 validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies for ResNet-101 on ImageNet for $H=32$.
+Figure 9: Validation accuracy, top-5 validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies for ResNet-101 on ImageNet for $H=16$.
+Figure 10: Validation accuracy, top-5 validation accuracy and local batch sizes curves of Local SHB with adaptive batch size strategies for ResNet-101 on ImageNet for $H=4$.